### v2

# Fast constrained sampling in pre-trained diffusion models

Abstract
Large denoising diffusion models, such as Stable Diffusion, have been trained on billions of image-caption pairs to perform text-conditioned image generation. As a byproduct of this training, these models have acquired general knowledge about image statistics, which can be useful for other inference tasks. However, when confronted with sampling an image under new constraints, e.g. generating the missing parts of an image, using large pre-trained text-to-image diffusion models is inefficient and often unreliable. Previous approaches either utilize backpropagation, making them significantly slower and more memory-demanding than text-to-image inference, or only enforce the constraint locally, failing to capture critical long-range correlations. In this work, we propose an algorithm that enables fast and high-quality generation under arbitrary constraints. We observe that, during inference, we can interchange between gradient updates computed on the noisy image and updates computed on the final, clean image. This allows us to employ a numerical approximation to expensive gradient computations, incurring significant speed-ups in inference. Our approach produces results that rival or surpass the state-of-the-art training-free inference approaches while requiring a fraction of the time. We demonstrate the effectiveness of our algorithm under both linear and non-linear constraints. An implementation is provided at this link.
Diffusion models, conditional sampling
1 Introduction
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
Constraint PSLD - 8min FreeDoM - 1min MPGD - 18s Ours - 15s Ground Truth
(Rout et al., 2023) (Yu et al., 2023) (He et al., 2024)
Figure 1: When tasked with completing the missing half of an image, previous methods are slow and fail to capture the important long-range dependencies between image pixels. The proposed algorithm generates a plausible image at a fraction of the time.
The development of large text-to-image models (Nichol et al., 2022; Rombach et al., 2022; Ramesh et al., 2022; Saharia et al., 2022b) has made denoising diffusion (Sohl-Dickstein et al., 2015; Ho et al., 2020) the go-to approach for capturing complex data distributions in high-dimensional spaces, such as images. By training on billions of text-image pairs, these models have acquired general knowledge about the image space, beyond text-to-image generation. This knowledge has been useful in quickly adapting to new conditions (Zhang et al., 2023) as well as utilizing model features to solve downstream image tasks (Tang et al., 2023).
The simplest way to utilize this knowledge is by fine-tuning the model. However, fine-tuning may require non-trivial computational resources and thus, previous works have focused on developing algorithms for conditional generation using only the pre-trained model (Chung et al., 2023; Rout et al., 2023; Chung et al., 2024; Yu et al., 2023; He et al., 2024). These methods modify the diffusion sampling process by computing additional gradient terms that move the sample towards the condition while denoising. When these gradients are computed using backpropagation through the model weights, there is a significant increase in inference time. On the other hand, when attempting to save computation by not propagating the condition information through the model, the generated image fails to capture the necessary long-range correlations.
As an example, in Figure 1, we use different algorithms to fill in the missing half of an image. Methods that backpropagate through the denoiser model (LDPS, PSLD (Rout et al., 2023), FreeDoM (Yu et al., 2023)) require significantly more time to run and do not consistently produce realistic results. Algorithms that do not compute gradients through the denoiser (MPGD (He et al., 2024)) fail to propagate the condition to distant pixels.
With the shortcomings of existing approaches in mind, we pose the following question: Can we avoid backpropagation through the model weights while at the same time maintaining long-range consistency when updating the sample with the conditioning signal? We observe that when a constraint increases the intensity of pixels in the predicted clean image, these pixel values are also expected to increase in the noisy image. From the diffusion point of view, the diffusion latent ${\bm{x}}_{t-1}$ is a linear combination of the current noisy image ${\bm{x}}_{t}$ and the predicted final image $\hat{{\bm{x}}}_{0}$ (Song et al., 2020), therefore any change in ${\bm{x}}_{t}$ can be also performed by changes in $\hat{{\bm{x}}}_{0}$.
This observation tells us that changes computed in the noisy and clean images can be interchanged. We find that this can be exploited to devise a new training-free conditional sampling algorithm that employs a numerical approximation for the expensive gradient computations. This allows us to speed up inference time while maintaining long-range dependencies in the image (Figure1). We take inspiration from Newton’s method and pose the proposed algorithm as an approximation to Newton steps that avoids computing the Jacobian inverse by using the Jacobian matrix directly.
We test the effectiveness of our proposed approach by generating images under both linear and non-linear constraints. We first show that our approach matches the results of state-of-the-art methods on free-form inpainting and 8×8$\times$8 × super-resolution at a fraction of the inference time. We then demonstrate how existing methods fail at inpainting large regions (25% of the image), while our algorithm obtains results closer to explicitly fine-tuning the diffusion model on inpainting. Finally, we apply our algorithm on non-linear constraints to perform style-guided and mask-guided generation, where we show that the proposed method synthesizes images that satisfy the constraints better than existing approaches.
2 Background
2.1 Denoising Diffusion Models
Denoising diffusion models, proposed for image generation by Ho et al. (2020), have now been widely adopted mainly due to their exceptional ability to synthesize diverse and high-quality samples. The original formulation views the training and inference process as a hierarchical latent variable model ${\bm{x}}_{T}\rightarrow{\bm{x}}_{T-1}\rightarrow\dots\rightarrow{\bm{x}}_{1}\rightarrow{\bm{x}}_{0}$, where the final latent is distributed normally ${\bm{x}}_{T}~{}\sim N(\mathbf{0},{\bm{I}})$ and $p({\bm{x}}_{0})$ represents the data distribution. After choosing a noise schedule $a_{t}$ that defines the forward transitions ${\bm{x}}_{t}\rightarrow{\bm{x}}_{t+1}$ that corrupt the data with Gaussian noise, usually centered at $\sqrt{\frac{a_{t+1}}{a_{t}}}{\bm{x}}_{t}$ and with variance $(1-\frac{a_{t+1}}{a_{t}})$, the model is trained to reverse each individual step in the diffusion process by predicting the noise added to the clean sample. In this formulation, the predicted noise is shown to approximate the score function $\nabla_{{\bm{x}}_{t}}\log p_{t}({\bm{x}}_{t})$ (Song et al., 2021) of the diffusion latent variables.
Further iterations of denoising diffusion introduced classifier guidance (Dhariwal & Nichol, 2021), which adapted a pre-trained unconditional diffusion model for conditional sampling by training an additional classifier $p({\bm{y}}|{\bm{x}}_{t})$. During inference, each reverse step is also perturbed by the gradient $\nabla_{{\bm{x}}_{t}}\log p({\bm{y}}|{\bm{x}}_{t})$ which guides the diffusion latent ${\bm{x}}_{t}$ towards regions that also satisfy the condition ${\bm{y}}$. Classifier-free guidance (Ho & Salimans, 2022) eliminated the need for training an additional classifier to perform conditional generation by incorporating it into the base diffusion model training process.
Latent diffusion models (LDMs) proposed Rombach et al. (2022), reduced the complexity of training large diffusion models by training the diffusion in the latent space of a pre-trained autoencoder. By utilizing an encoder $\mathcal{E}$ and decoder $\mathcal{D}$ that accurately reconstruct images ${\bm{x}}_{0}\approx\mathcal{D}(\mathcal{E}({\bm{x}}_{0}))$, the diffusion process can be made more efficient as the learned redundant information in an image is left for the decoder to reconstruct. Most large text-to-image diffusion models, such as Stable Diffusion (Podell et al., 2024), are based on this encoder-decoder approach to reduce complexity.
2.2 Gradient descent steps for constrained sampling
Previous works have studied whether large pre-trained diffusion models, which required significant investment to train, can be directly used for inference under novel conditions without additional tuning for each different constraint (Chung et al., 2023). The most typical problem formulation is denoising the sequence ${\bm{x}}_{T},{\bm{x}}_{T-1},...,{\bm{x}}_{1},{\bm{x}}_{0}$ under the linear constraint on the final signal in the form ${\bm{A}}{\bm{x}}_{0}={\bm{y}}$, or using a relaxed version that requires minimization of $||{\bm{A}}{\bm{x}}_{0}-{\bm{y}}||_{2}^{2}$, possibly as part of the likelihood function $p({\bm{y}}|{\bm{x}}_{0})={\cal N}({\bm{y}};{\bm{A}}{\bm{x}}_{0},\sigma^{2}{\bm{I}})$.
In contrast to classifier guidance, which explicitly trained a separate model for the likelihood $p({\bm{y}}\mid{\bm{x}}_{t})$, the linear constraint only applies to the final, noise-free image ${\bm{x}}_{0}$. Thus, existing methods rely on Tweedie’s formula (Efron, 2011), by which denoising diffusion models approximating $\nabla_{{\bm{x}}_{t}}\log p_{t}({\bm{x}}_{t})$ can be used to express the expected value of ${\bm{x}}_{0}$, denoted as $\hat{{\bm{x}}}_{0}$. Including the constraint as if an additional observed variable ${\bm{y}}$ was generated requires the addition of the term $\nabla_{{\bm{x}}_{t}}\log p({\bm{y}}\mid{\bm{x}}_{t})$ to every diffusion sampling step, and most previous methods focused on different approximations of $p({\bm{y}}\mid{\bm{x}}_{t})$ using the estimated $\hat{{\bm{x}}}_{0}$.
Regardless of how the constraint gradient is applied, the regular denoising diffusion steps are altered so that at each $t$, the generated latent ${\bm{x}}_{t}$ is moved in the direction reducing the cost
$C({\bm{x}}_{t})=({\bm{A}}\hat{{\bm{x}}}_{0}({\bm{x}}_{t})-{\bm{y}})^{T}({\bm{A}}\hat{{\bm{x}}}_{0}({\bm{x}}_{t})-{\bm{y}}).$ (1)
For example, in the case of inpainting missing pixels, the matrix ${\bm{A}}$ extracts the subsection of the known pixels in image ${\bm{x}}_{0}$ to be compared with a given target ${\bm{y}}$. The estimated expected value of ${\bm{x}}_{0}$ at the end of the chain is provided by the diffusion model as a nonlinear function $\hat{{\bm{x}}}_{0}({\bm{x}}_{t})$ learned by the denoiser network during training. Typically, these moves are gradient descent moves, i.e. moves of ${\bm{x}}_{t}$ in the direction of $-{\bm{e}}_{t}$ where
$-{\bm{e}}_{t}=-\nabla_{{\bm{x}}_{t}}C({\bm{x}}_{t})=-{\bm{J}}^{T}{\bm{A}}^{T}({\bm{A}}\hat{{\bm{x}}}_{0}-{\bm{y}})=-{\bm{J}}^{T}{\bm{e}},$ (2)
${\bm{J}}=\nabla_{{\bm{x}}_{t}}\hat{{\bm{x}}}_{0}({\bm{x}}_{t}),\quad{\bm{e}}={\bm{A}}^{T}({\bm{A}}\hat{{\bm{x}}}_{0}-{\bm{y}}).$ (3)
Note that we name the error signal ${\bm{e}}$ as the (negative) direction of the gradient w.r.t. ${\bm{x}}_{0}$ itself, and the ${\bm{e}}_{t}$ is the matching move in ${\bm{x}}_{t}$.
The matrix ${\bm{A}}^{T}$ inverses the operation of ${\bm{A}}$ and can usually be computed for each task a-priori. However, computation of the full Jacobian ${\bm{J}}$ is impractical, so the gradient is instead computed using backpropagation through $C({\bm{x}}_{t})$. In (Chung et al., 2023) and (Chung et al., 2024) for example, gradient steps of this form are applied to ${\bm{x}}_{t}$ at each step $t$ of generation before moving on to the next stage. As optimization of ${\bm{x}}_{t}$ might reduce the total noise in the image below what the denoising at $t-1$ was trained for, the gradient steps moving ${\bm{x}}_{t}$ towards optimizing $C({\bm{x}}_{t})$ can be combined with adding additional noise, which could also be seen as a form of stochastic averaging, as done by Yu et al. (2023).
3 Method
3.1 Approximate Newton steps
We start by observing that a Newton optimization step, instead of moving ${\bm{x}}_{t}$ in the gradient descent direction ${\bm{J}}^{T}{\bm{e}}$, moves it in the direction ${\bm{J}}^{-1}{\bm{e}}$. We demonstrate that on a more general form of the cost
$C({\bm{x}}_{t})=(f(\hat{{\bm{x}}}_{0}({\bm{x}}_{t}))-{\bm{y}})^{T}(f(\hat{{\bm{x}}}_{0}({\bm{x}}_{t}))-{\bm{y}}),$ (4)
for some target ${\bm{y}}$ to be matched with projection function $f$, which in the linear case is $f({\bm{x}})={\bm{A}}{\bm{x}}$. The Gauss-Newton optimization would first approximate
$f(\hat{{\bm{x}}}_{0}({\bm{x}}_{t}-{\bm{e}}_{t}))\approx f(\hat{{\bm{x}}}_{0}({\bm{x}}_{t}))-{\bm{J}}_{f}{\bm{J}}{\bm{e}}_{t},$ (5)
where ${\bm{J}}_{f}$ is the Jacobian of $f$. We then rewrite the cost of move $-{\bm{e}}_{t}$ as
$C({\bm{x}}_{t}-{\bm{e}}_{t})=$ $(f(\hat{{\bm{x}}}_{0}({\bm{x}}_{t}))-{\bm{J}}_{f}{\bm{J}}{\bm{e}}_{t}-{\bm{y}})^{T}$
$(f(\hat{{\bm{x}}}_{0}({\bm{x}}_{t}))-{\bm{J}}_{f}{\bm{J}}{\bm{e}}_{t}-{\bm{y}}).$ (6)
The cost is minimized by setting $\nabla_{{\bm{e}}_{t}}C=0$ to get the system
${\bm{J}}^{T}{\bm{J}}_{f}^{T}(f(\hat{{\bm{x}}}_{0}({\bm{x}}_{t}))-{\bm{y}})={\bm{J}}^{T}{\bm{J}}_{f}^{T}{\bm{J}}_{f}{\bm{J}}{\bm{e}}_{t}$
$\Leftrightarrow{\bm{e}}_{t}={\bm{J}}^{-1}{\bm{J}}_{f}^{-1}(f(\hat{{\bm{x}}}_{0}({\bm{x}}_{t}))-{\bm{y}})$ (7)
${\bm{e}}_{t}={\bm{J}}^{-1}{\bm{e}},\quad{\bm{e}}={\bm{J}}_{f}^{-1}(f(\hat{{\bm{x}}}_{0}({\bm{x}}_{t}))-{\bm{y}}),$ (8)
assuming the inverses exist. In the case of $f({\bm{x}})={\bm{A}}{\bm{x}}$ for inpainting and superresolution tasks, where ${\bm{y}}$ is lower-dimensional, the inverse of ${\bm{J}}_{f}$, obviously is not defined, but we can use the pseudoinverse. Similarly, for a non-linear $f$, we can compute ${\bm{e}}$ using numerical methods, e.g. backpropagation through $f$ when it is a neural network.
Therefore, ${\bm{e}}$ is the same in both gradient descent and Newton optimization, but the relationship to ${\bm{e}}_{t}$ is different:
$\text{GD}:\ {\bm{e}}_{t}={\bm{J}}^{T}{\bm{e}},\ \ \text{Newton}:\ {\bm{J}}{\bm{e}}_{t}={\bm{e}}$ (9)
The gradient descent step can be computed without directly evaluating the Jacobian by back-propagation on the scalar cost $C$. For the computation of ${\bm{J}}^{-1}{\bm{e}}$ on the other hand, we have no such method. However, we can efficiently compute the move
$\epsilon{\bm{e}}_{t}=\epsilon{\bm{J}}{\bm{e}}.$ (10)
We follow the Newton recipe suggesting multiplication of the error by ${\bm{J}}$ in Eq. (9) but multiply ${\bm{e}}$ by it instead of ${\bm{e}}_{t}$. Intuitively, this is justified by the fact that the mode of ${\bm{x}}_{t-1}$ in the next stage of denoising is a weighted sum of $\hat{{\bm{x}}_{0}}$ and ${\bm{x}}_{t}$, meaning that the changes in ${\bm{x}}_{t}$ can be equivalently represented by proportional moves in the same direction of either component. This is the consequence of the modeling assumption that in the posterior the stages are separated by independent Gaussian noise, and early inpaiting models simply assumed ${\bm{e}}_{t}={\bm{e}}$, missing on the effect of the image prior $p({\bm{x}}_{0})$, which is implicit in the learned function ${\bm{x}}_{0}({\bm{x}}_{t})$. In addition to that, in Appendix A.1 we provide an analysis of the convergence on the algorithm using inexact Newton methods (Dembo et al., 1982).
The direction of optimization we propose in Eq. (10) has at least two advantages over the usual gradient descent update. First, the direction $\epsilon{\bm{e}}_{t}$ can be computed numerically to save both on computation (2-fold) and memory (2.5-fold) compared to using backpropagation on the cost in Eq. (1). To derive that update, consider the function $f(s)=\hat{{\bm{x}}}_{0}({\bm{x}}_{t}-s{\bm{e}})$ where the variable $s$ is scalar. Its derivative at $s=0$ is
$\frac{df}{ds}=-{\bm{J}}{\bm{e}},$ (11)
and so the direction ${\bm{h}}$ can be approximated numerically using the approximate derivative of $f(s)$ at $s=0$
$f^{\prime}=\frac{1}{\delta}[\hat{{\bm{x}}}_{0}({\bm{x}}_{t}+\delta{\bm{e}})-\hat{{\bm{x}}}_{0}({\bm{x}}_{t})],$ (12)
${\bm{e}}_{t}=-\epsilon{\bm{J}}{\bm{e}}\approx\frac{\epsilon}{\delta}[\hat{{\bm{x}}}_{0}({\bm{x}}_{t}+\delta{\bm{e}})-\hat{{\bm{x}}}_{0}({\bm{x}}_{t})],$ (13)
requiring no backpropagation but instead two forward passes through the network: one to compute $\hat{{\bm{x}}}_{0}({\bm{x}}_{t})$ and another to compute $\hat{{\bm{x}}}_{0}$ for the ${\bm{x}}_{t}$ perturbed in the negative direction of the error vector.
The second advantage comes from the asymmetry of the Jacobian ${\bm{J}}$. The theoretically optimal ${\bm{x}}_{0}({\bm{x}}_{t})$ would have a symmetric Jacobian, as $\hat{{\bm{x}}}_{0}$ approximates the true expectation $E[{\bm{x}}_{0}|{\bm{x}}_{t}]=\frac{1}{\sqrt{\alpha_{t}}}[{\bm{x}}_{t}+{\bm{v}}_{t}\nabla_{{\bm{x}}_{t}}\log p_{t}({\bm{x}}_{t})]$, and the gradient of this is indeed symmetric:
$\nabla_{{\bm{x}}_{t}}E[{\bm{x}}_{0}|{\bm{x}}_{t}]=\frac{1}{\sqrt{\alpha_{t}}}[{\bm{I}}+{\bm{v}}_{t}\nabla^{2}\log p_{t}({\bm{x}}_{t})].$ (14)
This would render the gradient descent and the proposed update equivalent. In Section 3.2, we show that in practice, the trained denoiser models do not satisfy this ideal condition, making the two update directions noticeably different. Whether the denoising diffusion models are trained with score matching in mind (Song et al., 2021) or using the variational method of Ho et al. (2020), they do not directly optimize to match the real score $\nabla_{{\bm{x}}_{t}}\log p_{t}({\bm{x}}_{t})$ everywhere nor are they constrained to produce symmetric Jacobians. We hypothesize that this discrepancy in the two update directions is one of the reasons why our optimization of ${\bm{x}}_{t}$ may be more suitable for some applications, as demonstrated in Section 4.
Lastly, the proposed update of Eq. (10) only requires us to provide the direction of the error ${\bm{e}}$, which points locally towards a lower cost for our constraint on ${\bm{x}}_{0}$. As already discussed, for linear constraints such as inpainting, this direction can be obtained in closed form using the inverse of the operator expressed by ${\bm{A}}$. For non-linear constraints, such as style-guidance, ${\bm{e}}$ can be computed using backpropagation through the differentiable non-linear constraint model. Note that this backpropagation operation does not backpropagate through the denoiser itself, but only through the constraint w.r.t. the estimated $\hat{{\bm{x}}}_{0}$. A general version of the proposed algorithm for linear and non-linear constraints is presented in Algorithm 1. We choose DDIM as our sampling algorithm (Song et al., 2020). Applying other sampling methods should be intuitive by interleaving the diffusion and our gradient updates on ${\bm{x}}_{t}$.
Algorithm 1 The proposed algorithm for constrained sampling using a pre-trained diffusion model.
1:  Input: Pre-trained diffusion model $\hat{{\bm{x}}}_{0}({\bm{x}}_{t})$, constraint $C({\bm{x}}_{0},{\bm{y}})=\lVert f(\hat{{\bm{x}}}_{0}({\bm{x}}_{t})-y\rVert_{2}^{2}$, condition ${\bm{y}}$, gradient step size $\delta$, iterations $K$, learning rate $\lambda$, diffusion step size $s$ and schedule parameters $\alpha_{i},\beta_{i}$
2:  ${\bm{x}}_{T}\sim N(\textbf{0},{\bm{I}})$
3:  for $t\in\{T,T-s,T-2s,\dots,s\}$ do
4:     for $i\in\{1,2,\dots,K\}$ do
5:        ${\bm{e}}={\bm{J}}_{f}^{-1}(f(\hat{{\bm{x}}}_{0}({\bm{x}}_{t}))-{\bm{y}})$Task-dependent}
6:        ${\bm{e}}_{t}=[\hat{{\bm{x}}}_{0}({\bm{x}}_{t}+\delta{\bm{e}})-\hat{{\bm{x}}}_{0}({\bm{x}}_{t})]/\delta$
7:        ${\bm{x}}_{t}={\bm{x}}_{t}+\lambda{\bm{e}}_{t}$
8:     end for
9:     ${\bm{z}}_{t}\sim N(\textbf{0},{\bm{I}})$
10:     $\mathbf{\epsilon}_{t}=\frac{1}{\sqrt{1-\alpha_{t-s}}}{\bm{x}}_{t}-\frac{\sqrt{\alpha_{t-s}}}{\sqrt{1-\alpha_{t-s}}}\hat{{\bm{x}}}_{0}({\bm{x}}_{t})$
11:     ${\bm{x}}_{t-s}=\sqrt{\alpha_{t-s}}\hat{{\bm{x}}}_{0}({\bm{x}}_{t})+\sqrt{1-\alpha_{t-s}}\mathbf{\epsilon}_{t}+\beta_{t-s}{\bm{z}}_{t}$
12:  end for
13:  Return: ${\bm{x}}_{0}$
3.2 Symmetry of the Jacobian and difference in updates
Figure 2: Sample pairs $(i,j),(k,l)$ of the denoiser Jacobian $\nabla_{{\bm{x}}_{t}}\hat{{\bm{x}}_{0}}=\nabla_{{\bm{x}}_{t}}E[{\bm{x}}_{0}|{\bm{x}}_{t}]$ for different timesteps $t$. We observe that the Jacobian is not symmetric, justifying the difference between the proposed update steps and the previously used gradient descent steps.
Although, theoretically, the Jacobian of the denoiser should be symmetric (Eq. 14), the trained diffusion model does not exactly match the real score and can yield non-symmetric Jacobians. In Figure 2, we verify this claim with a simple experiment; we select a random image from the ImageNet (Deng et al., 2009) validation set and employ the Stable Diffusion 1.5 model (Podell et al., 2024) to denoise at three different noise levels $t=\{900,700,400\}$. We encode the image using the VAE encoder, scale it and add appropriate noise to get the intermediate diffusion latent for each timestep. We then give the noisy image to the denoiser network and compute the gradients $\partial\hat{{\bm{x}}}_{0}^{k,l}/\partial{\bm{x}}_{t}^{i,j}$ and $\partial\hat{{\bm{x}}}_{0}^{i,j}/\partial{\bm{x}}_{t}^{k,l}$ for randomly chosen pixels $(i,j),(k,l)$ using backpropagation. When we plot the gradients and compute the correlation coefficient $r$ we observe that the values deviate from $y=x$, which would indicate a symmetric Jacobian.
Figure 3: We present a simple experiment to compare the proposed gradient update ${\bm{J}}{\bm{e}}$ with that of previous methods ${\bm{J}}^{T}{\bm{e}}$. The goal is to update a noisy version of the image shown in (a) at $t=800$, such that the diffusion model’s prediction of the final image includes a red square in the middle, shown in (b). The results of gradient descent in (c) and our gradient update in (d) demonstrate that updates performed by previous methods lead to a considerably different result than the update we propose in this paper.
Having shown that the Jacobian is not symmetric, we expect to find differences between the gradient updates of gradient descent (Eq. 3) and our proposed update (Eq. 10). To highlight these differences, we set up a toy experiment where we update an identical initial image with the two different directions, under the same condition. The experiment is showcased in Figure 3 and again utilizes the Stable Diffusion 1.5 model. We create a synthetic black-and-white grid image (Figure 3 (a)), which we blur, deterministically encode, scale and noise to get a diffusion latent at $t=800$. We then set up a constraint where we add a red square to the center of the original image (Figure 3 (b)). The brown region in the constraint indicates masked-out values, or equivalently, where we will not be applying a constraint. Instead of decoding the image and applying the constraint in image space, we also encode the constraint image and apply it directly in the Stable Diffusion latent space.
We run 5 gradient updates with the same learning rate of 1 for each and demonstrate the final predicted $\hat{{\bm{x}}}_{0}$ for the ${\bm{J}}^{T}{\bm{e}}$ update of (3) (Figure 3 (c)) and the proposed ${\bm{J}}{\bm{e}}$ (10) (Figure 3 (d)). The final result shows how the model intends to change the entire image when asked to add a red square in the middle. We find that the resulting images differ substantially, with the proposed direction producing a more coherent image that tries to copy the newly introduced texture to the correct locations, i.e. the intersections of the lines. Although this is an empirical observation, we hypothesize that the two different directions can have vastly different effects on the image. In the Appendix A.2, we further visualize differences between ${\bm{J}}^{T}{\bm{e}}$ and ${\bm{J}}{\bm{e}}$, and in C.5, we repeat the previous experiment over multiple seeds.
4 Experiments
4.1 Linear Constraints
Figure 4: Comparison between our method and existing algorithms on free-form inpainting and 8×8$\times$8 × super-resolution. We directly use the images and results from (Chung et al., 2024) since there is no code available to replicate their method.
Inpaint (Free-form) SR (×$\times 8$)
Method PSNR ↑ LPIPS ↓ FID ↓ PSNR ↑ LPIPS ↓ FID ↓ Time
P2L (Chung et al., 2024) 21.99 0.229 32.82 23.38 0.386 51.81 30min*
LDPS 21.54 0.332 46.72 23.21 0.475 61.09 6min
PSLD (Rout et al., 2023) 20.92 0.251 40.57 23.17 0.471 60.81 8min
Ours 21.73 0.258 19.39 24.26 0.455 60.99 15s/1min
Table 1: Quantitative evaluation (PSNR, LPIPS, FID) on free-form inpainting and 8×8$\times$8 × superresolution on ImageNet. Our algorithm is competitive with the results of P2L which requires at least 30×30$\times$30 × more time for a single sample.
Figure 5: Qualitative evaluation of large area (box) inpainting on ImageNet. Our method provides results close to the fine-tuned inpainting model while requiring a fraction of the time to run per image.
We first verify our algorithm by generating images under linear constraints, which has been the main application of many previous algorithms (Chung et al., 2023; Rout et al., 2023; Chung et al., 2024). We follow the evaluation setting of Saharia et al. (2022a) and test our method on ImageNet (Deng et al., 2009), using the first 1000 images from the 10k validation set of the ctest10k split. As our evaluation metrics, we measure the PSNR and LPIPS (Zhang et al., 2018) between the generated and real images, as well as the FID (Heusel et al., 2017) between the real and generated sets. We use Stable Diffusion 1.4, which is pre-trained on the LAION (Schuhmann et al., 2022) text-image pair dataset. All experiments were run on NVIDIA RTX A5000 24GB GPUs.
Free-form Inpainting & Super-resolution
We follow the previously described experiment settings of free-form inpainting and 8×8$\times$8 × super-resolution. For free-form inpainting, the masks are randomly sampled and mask out 10-20% of the image pixels, as done by Saharia et al. (2022a). For inpainting with our method, we opt to directly operate on the Stable Diffusion latent given that Stable Diffusion VAE mostly compresses information locally. We apply the masking in pixel space, encode the masked image and inpaint with the unmasked VAE latents. We also apply a $3\times 3$ dilation kernel on the pixel mask before downsampling, masking out some extra pixels along the edge that we find the VAE fails to encode. For inpainting, we set the number of optimization steps $K=5$ over which we linearly decrease the learning rate $\lambda$ from 0.5 to 0.1. For super-resolution, we use $K=10$ and a constant $\lambda=0.1$. We perform only 20 DDIM (Song et al., 2020) steps for all experiments.
Regarding super-resolution, we cannot apply the constraint in the VAE latent space as image pixel downsampling does not correspond to downsampling VAE latents. To compute the error direction ${\bm{e}}$ we backpropagate the pixel-level constraint cost $({\bm{A}}\mathcal{D}({\bm{x}}_{t})-{\bm{y}})^{T}({\bm{A}}D({\bm{x}}_{t})-{\bm{y}})$ through the decoder network and utilize the resulting gradient w.r.t. the latent as the error direction. We note that by setting ${\bm{e}}=-\partial C(\mathcal{D}({\bm{x}}_{t}))/\partial{\bm{x}}_{t}$ we only require backpropagation through the decoder model, which is significantly less expensive than backpropagating through the denoiser network, as done by previous works. For both degradations, we also include additive white Gaussian noise with $\sigma_{\bm{y}}=0.05$.
We present the results on free-form inpainting and 8×8$\times$8 × super-resolution (SR) in Table 1. In inpainting, we find that our method generates better-aligned parts for the missing image regions, reflected in the significant improvement in FID, while obtaining similar PSNR and LPIPS scores to state-of-the-art methods. For super-resolution, although the improvements are smaller, we still find that we can compete with existing methods in both the quality and faithfulness of the generated images.
The main advantage we showcase is inference time; our approach achieves similar or better results to every other existing method while only requiring a fraction of the computation. When we do not perform backpropagation through the VAE decoder (inpainting), our method requires only 15 seconds. In super-resolution, where we perform more optimization steps and backpropagate, the time increases to 1 minute. In contrast, P2L (Chung et al., 2024), the best-performing method, requires 30 minutes per single image since it relies on additional backpropagation steps through the denoiser to infer the text condition. Since there is no publicly available implementation for P2L we report the results directly from the paper and estimate the inference time as 5×5$\times$5 × that of LDPS (Chung et al., 2023), which the authors mentioned as a reasonable expectation as it linearly scales with the number of text prompt optimization iterations.
Box Inpainting
In the previous experiments, we observed that our method performed better in inpainting, which required the model to infer more missing information in the given image than in super-resolution. We further push the model by asking it to inpaint a box that covers 25% of the input image. In the comparisons we also include FreeDoM (Yu et al., 2023), which, although not previously shown on inpainting, claims to be a fast training-free inference approach for any condition. We also include the fine-tuned SD-Inpaint model (Rombach et al., 2022), which required 500k additional training steps.
We present the results on box inpainting in Table 2 and Figure 5. Our approach outperforms all existing training-free methods in all metrics and is the closest to the fine-tuned SD-Inpaint model, which we consider an upper limit. Qualitatively, we find that LDPS and PSLD, which perform a lot of denoising steps, generate blurry parts that align with the average appearance over the rest of the image. FreeDoM, the faster baseline, although generating non-blurry parts, frequently fails to maintain consistency with the rest of the image. Our method achieves both high quality and consistency in a shorter time as all previous methods backpropagate through the denoiser weights. In Appendix C.1 we ablate the steps $K$ and learning rate $\lambda$ hyperparameters.
Inpaint (Box)
Method PSNR ↑ LPIPS ↓ FID ↓ Time
LDPS 17.52 0.42 76.32 6min
PSLD 17.30 0.38 74.02 8min
FreeDoM 16.18 0.42 55.68 1min
(Yu et al., 2023)
Ours 18.30 0.30 42.01 15s
SD-Inpaint 19.05 0.28 32.93 4s
Table 2: Quantitative evaluation (PSNR, LPIPS, FID) on large area (box) inpainting.
4.2 Style-guided generation
Our algorithm can be applied to non-linear differentiable constraints using backpropagation to approximate ${\bm{e}}$. Computing this gradient direction does not require backpropagation through the denoiser but through the separate, more lightweight ’critic’ network $f$. We demonstrate this with the non-linear constraint of style-guided generation.
The goal is to generate an image that simultaneously follows the style of a reference image ${\bm{x}}_{ref}$ and a given text prompt. Following previous works (Yu et al., 2023; He et al., 2024) which also perform style-guided generation, we do this by matching the statistics of CLIP (Radford et al., 2021) features between the reference and the generated images. We minimize the Frobenius norm between the Gram matrices (Gatys et al., 2016) of the 2nd layer CLIP features, making the cost to be minimized
$\|\text{Gram}(\text{CLIP}_{2}({\bm{x}}_{ref}))-\text{Gram}(\text{CLIP}_{2}(\hat{{\bm{x}}}_{0}))\|_{F}^{2}.$ (15)
Adhering to the evaluation of Yu et al (2023), we use 1000 random pairs of reference style images from WikiArt (Tan et al., 2019) and prompts from PartiPromtps (Yu et al., 2022). We measure the CLIP similarity between the generated images and the text prompts to evaluate the faithfulness to the text condition, and the final difference between the Gram matrices to evaluate the faithfulness to the style reference (Style Score). We use the CLIP ViT-B/16 model for guiding the style of the image and evaluating. We also repeat the experiment using the OpenCLIP ViT-B/32 model (Cherti et al., 2023) for guidance. We perform $K=5$ gradient updates for every denoising step, using a linearly decreasing learning rate $\lambda$ from 0.5 to 0.1 and classifier-free guidance (Ho & Salimans, 2022) $w=2$ and $w=5$ for the denoiser.
The results are presented in Table 3 and Figure 6. We find that our method achieves the overall best style score. Our approach is the best at minimizing the constraint which is also the the evaluation metric. Although in general, the lower the style score the more difficult it is to maintain high CLIP similarity with the prompt, we observe that our method still balances between the style and text, especially when we increase the guidance weight. We also used an OpenCLIP model to guide the style. With a CLIP model still as the evaluator, we achieve better style scores than the baselines that optimized directly with CLIP. Our proposed update reduces the target cost better without generating adversarial artifacts that trick the CLIP evaluation.
Regarding inference time, our method is the fastest out of all baselines, excluding the non-style-guided DDIM. The baseline of MPGD (He et al., 2024), which is comparable in inference speed, does not backpropagate the style loss through the denoiser network, but only modifies the $\hat{{\bm{x}}}_{0}$ estimation at every denoising step. As we demonstrate in the supplementary, MPGD completely fails to propagate the constraint to distant pixels when the conditioning affects a specific region in the image. This makes it unusable in some constrained sampling settings, such as box inpainting.
Figure 6: Style-guided generation. We show the results of guiding Stable Diffusion with a CLIP and an OpenCLIP model, using classifier-free guidance $w=5$. The images generated by our algorithm follow the reference style better than previous algorithms.
Method Style Score* ↓ CLIP ↑ Time
DDIM 761.0 31.61 9s
FreeDoM 498.8 30.14 80s
MPGD (He et al., 2024) 441.0 26.61 50s
Ours (w=2) 368.37 23.95 45s
Ours (w=5) 310.96 24.57 45s
Ours (OpenCLIP, w=5) 434.45 25.94 45s
Table 3: Quantitative evaluation of style generation. *The Style Score is what the gradient is optimizing.
4.3 Mask-guided generation
As an additional non-linear constraint, we guide the Stable Diffusion model with a separately trained face segmentation network. We employ an off-the-shelf model from huggingface and set the constraint $C$ to be the KL divergence between the per-pixel segmentation classes predicted for a reference image and the generated image. Using 100 images from the FFHQ (Karras et al., 2019) validation set, we run both our method and the MPGD (He et al., 2024) baseline. We chose MPGD because as the fastest baseline that works very well with ’dense’ constraints, i.e. constraints that are applied to all pixels.
The results in Table 4 show that with a similar compute budget, our method achieves both faithfulness (mIoU between generated and reference images) and image quality (CLIP-FID). We used CLIP-FID because it performs better than Inception-FID on a small set of images (Kynkäänniemi et al., 2023). MPGD with a high weight ($\rho$) ’burns in’ the segmentation mask, leading to non-realistic images, whereas a lower weight does not produce images faithful to the mask. We provide qualitative examples of the generated images and the masks used in the Appendix C.5.
Method mIoU ↑ CLIP-FID ↓ Time
DDIM 0.09 48.78 3s
MPGD ($\rho=1$) 0.47 77.11 1min
MPGD ($\rho=0.5$) 0.36 56.38 1min
Ours 0.42 59.79 1min
Table 4: Mask-conditioned generation using 100 FFHQ validation set images as reference, using Stable Diffusion with the prompt ’a headshot photo’.
5 Conclusion
In this work, we presented a new algorithm for fast inference in pre-trained diffusion models under constraints. Our approach exploits the relationship between the noisy and clean images during the denoising process to perform a gradient update that requires no expensive backpropagation operations through the diffusion model. Our method produces results comparable to the state-of-the-art with a fraction of the inference time. Our approach offers a practical method to sample from large pre-trained generative image models under any condition and has the potential to enable training-free downstream applications that rely on utilizing a strong image prior.
Impact Statement
This paper presents work whose goal is to advance the field of Machine Learning. We acknowledge that there are many potential societal consequences of our work. Beyond the positives, we highlight the ability to edit images with the intent to deceive and mislead as requiring careful consideration.
Acknowledgments
Part of this work was done during an internship at Microsoft Research, Redmond. This research was also partially supported by NSF grants IIS-2123920, IIS-2212046.
Appendix A Analysis of the proposed algorithm
A.1 Convergence
We resort to inexact Newton methods (Dembo et al., 1982) to analyze the convergence and validity of the proposed approach. The original Newton’s method aims to solve the system
${\bm{J}}{\bm{e}}_{t}={\bm{e}}.$ (16)
In cases where computing the inverse of the Jacobian, as is the case of the diffusion denoiser Jacobian, inexact Newton methods propose first finding an approximation ${\bm{e}}_{t}^{*}$ to the solution of Eq (16) and then performing the Newton step using the approximate solution and reiterating until convergence. For inexact Newton methods to converge to the correct solution, the residual must be strictly reduced at every step
${\bm{r}}={\bm{e}}-{\bm{J}}{\bm{e}}_{t}^{*},\quad\frac{\lVert{\bm{r}}\rVert_{2}}{\lVert{\bm{e}}\rVert_{2}}<\eta,\ \eta\in[0,1).$ (17)
If we substitute the proposed update ${\bm{e}}_{t}^{*}=\lambda{\bm{J}}{\bm{e}}$ (Eq. 10) we get the residual
${\bm{r}}={\bm{e}}-\lambda{\bm{J}}^{2}{\bm{e}}=({\bm{I}}-\lambda{\bm{J}}^{2}){\bm{e}}.$ (18)
For this residual to be strictly reduced at every iteration of our algorithm, we need
$\frac{\lVert{\bm{r}}\rVert_{2}}{\lVert{\bm{e}}\rVert_{2}}=\frac{\lVert({\bm{I}}-\lambda{\bm{J}}^{2}){\bm{e}}\rVert_{2}}{\lVert{\bm{e}}\rVert_{2}}\leq\frac{\lVert{\bm{I}}-\lambda{\bm{J}}^{2}\rVert_{2}\lVert{\bm{e}}\rVert_{2}}{\lVert{\bm{e}}\rVert_{2}}=\lVert{\bm{I}}-\lambda{\bm{J}}^{2}\rVert_{2}<\eta.$ (19)
Therefore, we need to show that the spectral norm of the matrix ${\bm{I}}-\lambda{\bm{J}}^{2}$ is smaller than 1 for a correct choice of learning rate $\lambda$.
If ${\bm{J}}$ was diagonalizable (or more generally normal) then we could directly estimate $\lVert{\bm{I}}-\lambda{\bm{J}}^{2}\rVert_{2}$ using the largest eigenvalue of ${\bm{J}}$, following the spectral theorem. While we cannot directly assume that ${\bm{J}}$ is normal, we know that ${\bm{J}}$ should be ’almost’ symmetric, as is the optimal denoiser solution shown in Eq. 14. This means that we could express ${\bm{J}}={\bm{S}}+{\bm{K}}$, where ${\bm{S}}=({\bm{J}}+{\bm{J}}^{T})/2$ is the symmetric component and ${\bm{K}}=({\bm{J}}-{\bm{J}}^{T})/2$ the skew-symmetric component, and expect $\lVert{\bm{K}}\rVert\ll\lVert{\bm{S}}\rVert$. In that case, we can approximate $\lVert{\bm{I}}-\lambda{\bm{J}}^{2}\rVert_{2}$ as
$\lVert{\bm{I}}-\lambda{\bm{J}}^{2}\rVert_{2}\approx\lvert 1-\lambda\max_{i}\mu_{i}^{2}\rvert$ (20)
where $\mu_{i}$ are the eigenvalues of ${\bm{S}}$.
To get an estimate of the magnitudes of the largest eigenvalues of ${\bm{S}}$, which are real, and of the eigenvalues of ${\bm{K}}$, which are imaginary, we use the Arnoldi iteration (Arnoldi, 1951). For the Arnoldi iteration we only require access to the matrix-vector products ${\bm{J}}{\bm{v}}$ and ${\bm{J}}^{T}{\bm{v}}$, which can be computed using the approximation of Eq. (12) and backpropagation respectively.
Running the Arnoldi iteration algorithm on ${\bm{J}}$, $({\bm{J}}+{\bm{J}}^{T})/2$ and $({\bm{J}}-{\bm{J}}^{T})/2$ we plot the magnitude of the largest eigenvalue of each matrix in Fig. 7 (a). We see that the contribution of the symmetric part of the matrix is the strongest, validating our assumption that the diffusion model’s Jacobian ${\bm{J}}$ is ’almost’ symmetric.
If we approximate the spectral norm of Eq.( 19) using the symmetric component of the Jacobian we see that for a correct choice of $\lambda$ we can ensure that $\lvert 1-\lambda\max_{i}\mu_{i}\rvert<1$. Now we revisit the example of Fig. 1 where we inpaint half of the image. Instead of using the empirical learning rate used in the paper (linearly decreasing and gradient normalized by the $\infty$-norm) we compute the largest eigenvalue of $({\bm{J}}+{\bm{J}}^{T})/2$, $\mu$, and select different $\lambda$ so that we satisfy or violate the bound provided by Eq. (20).
In Fig. 7 (b) we show that the error is not reduced for learning rate values that consistently violate the proposed bound. As expected, our algorithm consistently reduces the error at every iteration for a small enough learning rate, where the bound is always satisfied. When we use the empirical learning rate, shown in Fig. 7 (c), we see that the algorithm bounces between satisfying and not satisfying the computed bound and ends up at a lower error than the constant learning rate of Fig. 7 (b). We posit that the normalization by the $\infty$-norm helps the learning rate adjust the step size such that it substantially reduces the error while not leading to divergence.
Figure 7: Using the proposed analysis we visualize in (a) the largest eigenvalue of ${\bm{J}}$ as well as its symmetric and skew-symmetric components. The largest eigenvalue follows closely the largest eigenvalue of the symmetric part of the matrix. In (b) we demonstrate the convergence of our method for different learning rates. In (c) we show that our adaptive learning rate scheme initially oscillates around the theoretical convergence bound and eventually settles well-below it. These initial oscillations may be important for the high-quality results, as using lower learning rate did not attain similar-quality images.
A.2 Difference between ${\bm{J}}$ and ${\bm{J}}^{T}$
In the main text, we referred to the difference between ${\bm{J}}$ and ${\bm{J}}^{T}$ as a motivating factor for our approach and measured that difference by comparing elements $(i,j)$ and $(j,i)$ of the Jacobian matrix (Figure 2). In Figure 8 we further visualize the difference between using the Jacobian and its transpose by numerically computing how the Jacobian and the Jacobian transpose of the entire image would look for a single pixel, i.e. which parts of the image are affected by a change in a single pixel.
In Figure 8 we simplify the notation and denote as ${\bm{J}}^{(i)}$ the Jacobian for a pixel $i$, which we compute by summing the squares of the matrix columns that correspond to this pixel’s latent values. If our approach were equivalent to backpropagation, i.e. ${\bm{J}}={\bm{J}}^{T}$, then a change in a single pixel would have a similar effect on the rest of the image, up to some noise because of the finite difference approximation.
Contrary to the expectations, we find a significant difference between our proposed direction and backpropagation, with our approach having a better effect on retaining shapes and symmetry across the image. We see that in many cases, the model is trying to change ’corresponding’ parts of the image together, i.e. change both eyes or ears simultaneously, giving us a better understanding of the knowledge that the model has acquired regarding the image space in general.
Appendix B Proposed algorithm as modification of DDIM
To apply the proposed method we modified the DDIM sampling algorithm (Song et al., 2020). We provide a side-by-side comparison to show the difference between the original DDIM and the proposed algorithm. Extending our algorithm to other sampling methods should be intuitive by alternating between gradient updates from our algorithm and the diffusion updates computed with the diffusion sampling algorithm used.
Algorithm 2 DDIM sampling (Song et al., 2020) using a pre-trained diffusion model.
1:  Input: Pre-trained diffusion model $\hat{{\bm{x}}}_{0}({\bm{x}}_{t})$, diffusion step size $s$ and schedule parameters $\alpha_{i},\beta_{i}$
2:  ${\bm{x}}_{T}\sim N(\textbf{0},{\bm{I}})$
3:  for $t\in\{T,T-s,T-2s,\dots,s\}$ do
4:     ${\bm{z}}_{t}\sim N(\textbf{0},{\bm{I}})$
5:     $\mathbf{\epsilon}_{t}=\frac{1}{\sqrt{1-\alpha_{t-s}}}{\bm{x}}_{t}-\frac{\sqrt{\alpha_{t-s}}}{\sqrt{1-\alpha_{t-s}}}\hat{{\bm{x}}}_{0}({\bm{x}}_{t})$
6:     ${\bm{x}}_{t-s}=\sqrt{\alpha_{t-s}}\hat{{\bm{x}}}_{0}({\bm{x}}_{t})+\sqrt{1-\alpha_{t-s}}\mathbf{\epsilon}_{t}+\beta_{t-s}{\bm{z}}_{t}$
7:  end for
8:  Return: ${\bm{x}}_{0}$
Algorithm 3 The proposed algorithm for constrained sampling using a pre-trained diffusion model.
1:  Input: Pre-trained diffusion model $\hat{{\bm{x}}}_{0}({\bm{x}}_{t})$, constraint $C({\bm{x}}_{0},{\bm{y}})=\lVert f(\hat{{\bm{x}}}_{0}({\bm{x}}_{t})-y\rVert_{2}^{2}$, condition ${\bm{y}}$, gradient step size $\delta$, iterations $K$, learning rate $\lambda$, diffusion step size $s$ and schedule parameters $\alpha_{i},\beta_{i}$
2:  ${\bm{x}}_{T}\sim N(\textbf{0},{\bm{I}})$
3:  for $t\in\{T,T-s,T-2s,\dots,s\}$ do
4:     for $i\in\{1,2,\dots,K\}$ do
5:        ${\bm{e}}={\bm{J}}_{f}^{-1}(f(\hat{{\bm{x}}}_{0}({\bm{x}}_{t}))-{\bm{y}})$Task-dependent}
6:        ${\bm{e}}_{t}=[\hat{{\bm{x}}}_{0}({\bm{x}}_{t}+\delta{\bm{e}})-\hat{{\bm{x}}}_{0}({\bm{x}}_{t})]/\delta$
7:        ${\bm{x}}_{t}={\bm{x}}_{t}+\lambda{\bm{e}}_{t}$
8:     end for
9:     ${\bm{z}}_{t}\sim N(\textbf{0},{\bm{I}})$
10:     $\mathbf{\epsilon}_{t}=\frac{1}{\sqrt{1-\alpha_{t-s}}}{\bm{x}}_{t}-\frac{\sqrt{\alpha_{t-s}}}{\sqrt{1-\alpha_{t-s}}}\hat{{\bm{x}}}_{0}({\bm{x}}_{t})$
11:     ${\bm{x}}_{t-s}=\sqrt{\alpha_{t-s}}\hat{{\bm{x}}}_{0}({\bm{x}}_{t})+\sqrt{1-\alpha_{t-s}}\mathbf{\epsilon}_{t}+\beta_{t-s}{\bm{z}}_{t}$
12:  end for
13:  Return: ${\bm{x}}_{0}$
Figure 8: We visualize the difference between ${\bm{J}}$ and ${\bm{J}}^{T}$ by computing how the matrix wants to change the entire image would look for a perturbation of a single pixel. By ${\bm{J}}^{(i)}$ we denote this exact change, which corresponds to the sum of 4 columns in the matrix (4 latent channels per-pixel). Our proposed update direction better captures longer-range dependencies by better maintaining shapes. Even though we use a numerical approximation, the proposed direction is sharper in regions, like the outline of the cat.
Appendix C Additional Results
C.1 Number of steps and learning rate ablation
We repeat the ImageNet box inpainting experiments with fewer optimization iterations and a higher learning rate. We show qualitative results in Figure 9, where we observe that running fewer optimization steps gives blurrier results, which is expected as the known regions of the image also seem to not have converged to the given values. Using a higher learning rate leads to the model sometimes ’overshooting’ by inpainting the missing regions with realistic-looking parts that do not necessarily fit the rest of the image. The quantitative results are presented in Table 5.
Method PSNR ↑ LPIPS ↓ FID ↓
Ours (n=5, lr=0.5) 18.30 0.30 42.01
Ours (n=2, lr=0.5) 18.01 0.39 68.75
Ours (n=5, lr=1.0) 17.48 0.32 47.20
Table 5: Ablation of the number of steps and learning rate on box inpainting. When using fewer steps, the algorithm does not sufficiently converge. When using a very high learning rate, the approach seems to overshoot to non-realistic inpainted parts.
Figure 9: Examples of box inpainting when using different optimization steps and learning rates.
C.2 Non-differentiable constraints
There is no restriction on defining the constraint $C$ as long as we can get the direction ${\bm{e}}$ towards which we want to push the image ${\bm{x}}_{0}$. In the non-linear constraint case, we resorted to using the gradient descent direction ${\bm{J}}_{f}^{T}(f(\hat{{\bm{x}}}_{0}({\bm{x}}_{t}))-{\bm{y}})$ to avoid computing the inverse Jacobian of $f$. In theory, any direction ${\bm{e}}$ that locally minimizes the constraint can be used with the proposed algorithm.
As a toy example, we generate images with pixel values quantized to be either ’on’ or ’off’ (-1 or 1). The constraint first measures whether a pixel value is positive or negative and then sets the error direction ${\bm{e}}$ to $|1-x|$ or $|-1-x|$ for every pixel accordingly. This is a non-differentiable constraint for which we can easily compute a local gradient that reduces the cost $C$. Using the prompt ’a photo of a cat’, we generate quantized images as shown in Figure 10.
Figure 10: We apply a non-differentiable constraint to generate quantized images. Using the prompt ’a photo of a cat’ we generated binary images of cats.
C.3 Inference visualizations
In Figures 11, 12, and 13, we visualize the intermediate steps of the proposed algorithm for the inpainting, super-resolution and style-guided generation tasks respectively. Our method quickly converges to a plausible image and then further refines it to better satisfy the constraint over the diffusion timesteps. For style-guided generation, we see that the structure of the image is defined in the first few initial steps before the specific style provided is applied.
Figure 11: Visualization of the intermediate inference steps for the inpainting task.
Figure 12: Visualization of the intermediate inference steps for the super-resolution task.
Figure 13: Visualization of the intermediate inference steps for the style-guided generation task.
C.4 Effect of convergence speed on final images
We ask the question of how does convergence speed affect the quality of the generated images? Previous works found that applying a high weight on the constraint led to unwanted artifacts in the generated images (Chung et al., 2023). We hypothesize that, apart from artifacts in the gradient, ’over-optimizing’ for the condition at a given timestep can affect the generation quality. In practice, if we push the initial ${\bm{x}}_{t}$ too far from the inputs the denoiser network is expecting, either with a high weight on the gradient update or by performing too many updates, we should be seeing non-realistic images in the output.
We investigate this by running the same inpainting experiment with 5 optimization steps per timestep (Figure 14) and 20 optimization steps (Figure 15). Although we expected to see a difference in the final generated images, we find that both converge to similar quality results. Our proposed optimization steps at a single timestep consider the Jacobian of the denoiser model, which we find acts as ’regularization’ and makes it difficult to produce $\mathbf{x}_{t}$ inputs that satisfy the condition ’early’. Even when running the optimization for more steps at a single $\mathbf{x}_{t}$, we see that although the sample converges faster to the desired condition, the denoiser is still able to continue the diffusion process of $\mathbf{x}_{t}$.
Figure 14: Convergence for the inpainting task when using 5 optimization steps $K$.
Figure 15: Convergence for the inpainting task when using 20 optimization steps $K$.
C.5 More Qualitative Results
In Figure 4 we provided qualitative results on free-form inpainting and super-resolution. In inpainting, our model consistently performs as well as the slowest baseline, P2L. For super-resolution, P2L which also infers a prompt seems to generate better-fitting textures for the images. We hypothesize that by also inferring a prompt the high-frequency detail generation is better-guided in the super-resolution task. In contrast, in inpainting, the non-masked pixels contain enough information about the textures that need to be placed around the image.
In Figure 16 we showcase additional results on the box inpainting task. MPGD (He et al., 2024), which does not backpropagate the constraint error through the diffusion model completely fails at inpainting the missing region. We attribute that to the minimal ability to influence pixels that are ’far’ from the constraint at lower noise levels without probing the model weights. The Naive algorithm replaces the known pixels in the estimated final image at every denoising iteration.
Figure 16: Box inpainting examples for all methods. Naive replaces the pixels with their true values + noise during inference.
In Figure 17 we present additional results on style-guided text-to-image generation for a single prompt. Qualitatively, we see that the style of the images generated with our algorithm better matches the style of the reference image, even when using a different model to define style (OpenCLIP). In Figure 18 we show images generated with different styles and text prompts. Here, we show how increasing the classifier-free guidance weight $w$ (Ho & Salimans, 2022) controls the influence of the text prompt on the final generated image.
Figure 17: Examples of style-guided text-to-image generation for a single prompt and multiple styles.
Figure 18: Examples of style-guided text-to-image generation for multiple prompts and styles.
In Figure 19 we provide qualitative results of our segmentation mask-guided generation experiment. We see that the baseline of MPGD either over-satisfies the constraint by burning in artifacts ($\rho=1$) or fails to generate images that adhere to the constraint ($\rho=0.5$). Our approach generates the most realistic images while also getting the mask prediction to match to the reference image. Since we are using Stable Diffusion, we used the text prompt ’a headshot photo’ to constrain the set of possible generated images. As expected, the results are not on par with an FFHQ-specific model, but we still find our algorithm able to synthesize good images even in this difficult case.
Figure 19: Examples of segmentation-guided generation using our method and MPGD.
In Figure 20 we include more samples from the experiment in Section 3.2 in the main paper. We observe that the proposed gradient updates regularly give more consistent outputs that repeat textures across the image. We attribute this to the difference between ${\bm{J}}$ and ${\bm{J}}^{T}$ we discussed in A.2.
Figure 20: Comparison between gradient descent and the proposed gradient. The setup is the same one as in Section 3.2.

### v3

@@
 generating the missing parts of an image, using large pre-trained text-to-image diffusion models is inefficient and often unreliable.
-Previous approaches either utilize backpropagation, making them significantly slower and more memory-demanding than text-to-image inference, or only enforce the constraint locally, failing to capture critical long-range correlations.
-In this work, we propose an algorithm that enables fast and high-quality generation under arbitrary constraints.
-We observe that, during inference, we can interchange between gradient updates computed on the noisy image and updates computed on the final, clean image.
-This allows us to employ a numerical approximation to expensive gradient computations, incurring significant speed-ups in inference.
-Our approach produces results that rival or surpass the state-of-the-art training-free inference approaches while requiring a fraction of the time.
-We demonstrate the effectiveness of our algorithm under both linear and non-linear constraints.
-An implementation is provided at this link.
-Diffusion models, conditional sampling
+Previous approaches either utilized backpropagation through the denoiser network, making them significantly slower and more memory-demanding than simple text-to-image generation, or only enforced the constraint locally, failing to capture critical long-range correlations in the sampled image.
+In this work, we propose an algorithm that enables fast, high-quality generation under arbitrary constraints.
+We show that in denoising diffusion models, we can employ an approximation to Newton’s optimization method that allows us to speed up inference and avoid the expensive backpropagation operations.
+Our approach produces results that rival or surpass the state-of-the-art training-free inference methods while requiring a fraction of the time.
+We demonstrate the effectiveness of our algorithm under both linear (inpainting, super-resolution) and non-linear (style-guided generation) constraints.
+An implementation is provided at this GitHub repository.
 1 Introduction
-Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
-Constraint PSLD - 8min FreeDoM - 1min MPGD - 18s Ours - 15s Ground Truth
-(Rout et al., 2023) (Yu et al., 2023) (He et al., 2024)
-Figure 1: When tasked with completing the missing half of an image, previous methods are slow and fail to capture the important long-range dependencies between image pixels.
-The proposed algorithm generates a plausible image at a fraction of the time.
-The development of large text-to-image models (Nichol et al., 2022; Rombach et al., 2022; Ramesh et al., 2022; Saharia et al., 2022b) has made denoising diffusion (Sohl-Dickstein et al., 2015; Ho et al., 2020) the go-to approach for capturing complex data distributions in high-dimensional spaces, such as images.
+The development of large text-to-image models [22, 27, 26, 30] has made denoising diffusion [32, 16] the go-to approach for capturing complex data distributions in high-dimensional spaces, such as images.
 By training on billions of text-image pairs, these models have acquired general knowledge about the image space, beyond text-to-image generation.
-This knowledge has been useful in quickly adapting to new conditions (Zhang et al., 2023) as well as utilizing model features to solve downstream image tasks (Tang et al., 2023).
-The simplest way to utilize this knowledge is by fine-tuning the model.
-However, fine-tuning may require non-trivial computational resources and thus, previous works have focused on developing algorithms for conditional generation using only the pre-trained model (Chung et al., 2023; Rout et al., 2023; Chung et al., 2024; Yu et al., 2023; He et al., 2024).
+This knowledge is useful in quickly adapting to new conditions [45, 40] and utilizing model features to solve downstream image tasks [38, 39].
+The simplest way to utilize this prior knowledge is by fine-tuning the model.
+However, fine-tuning may require non-trivial computational resources and thus, previous works have focused on developing algorithms for conditional generation using only the pre-trained model [5, 28, 6, 44, 13].
 These methods modify the diffusion sampling process by computing additional gradient terms that move the sample towards the condition while denoising.
@@
 As an example, in Figure 1, we use different algorithms to fill in the missing half of an image.
-Methods that backpropagate through the denoiser model (LDPS, PSLD (Rout et al., 2023), FreeDoM (Yu et al., 2023)) require significantly more time to run and do not consistently produce realistic results.
-Algorithms that do not compute gradients through the denoiser (MPGD (He et al., 2024)) fail to propagate the condition to distant pixels.
+Methods that backpropagate through the denoiser model (LDPS, PSLD [28], FreeDoM [44]) require significantly more time to run and do not consistently produce realistic results.
+Algorithms that do not compute gradients through the denoiser (MPGD [13]) fail to propagate the condition to distant pixels.
 With the shortcomings of existing approaches in mind, we pose the following question: Can we avoid backpropagation through the model weights while at the same time maintaining long-range consistency when updating the sample with the conditioning signal?
-We observe that when a constraint increases the intensity of pixels in the predicted clean image, these pixel values are also expected to increase in the noisy image.
-From the diffusion point of view, the diffusion latent ${\bm{x}}_{t-1}$ is a linear combination of the current noisy image ${\bm{x}}_{t}$ and the predicted final image $\hat{{\bm{x}}}_{0}$ (Song et al., 2020), therefore any change in ${\bm{x}}_{t}$ can be also performed by changes in $\hat{{\bm{x}}}_{0}$.
-This observation tells us that changes computed in the noisy and clean images can be interchanged.
-We find that this can be exploited to devise a new training-free conditional sampling algorithm that employs a numerical approximation for the expensive gradient computations.
-This allows us to speed up inference time while maintaining long-range dependencies in the image (Figure1).
-We take inspiration from Newton’s method and pose the proposed algorithm as an approximation to Newton steps that avoids computing the Jacobian inverse by using the Jacobian matrix directly.
-We test the effectiveness of our proposed approach by generating images under both linear and non-linear constraints.
-We first show that our approach matches the results of state-of-the-art methods on free-form inpainting and 8×8$\times$8 × super-resolution at a fraction of the inference time.
-We then demonstrate how existing methods fail at inpainting large regions (25% of the image), while our algorithm obtains results closer to explicitly fine-tuning the diffusion model on inpainting.
-Finally, we apply our algorithm on non-linear constraints to perform style-guided and mask-guided generation, where we show that the proposed method synthesizes images that satisfy the constraints better than existing approaches.
+We view the denoiser network as a function that removes all noise from an input image.
+In that context, the goal of constrained sampling is to find how the input should change such that the clean output satisfies the condition better.
+This involves the denoising function’s Jacobian matrix, which transforms a local change in the input to a change in outputs.
+We show that existing methods that backpropagate the constraint through the model employ the transpose of the Jacobian to invert this transformation, corresponding to the gradient descent direction.
+An alternative direction is given from the Newton method [24], but requires more computation as it involves finding the Jacobian inverse.
+However, the input and output of the denoiser are closely related; when the intensity of some pixels in the final, noise-free image is increased, the same pixels are also expected to brighten in the noisy image and vice versa.
+Therefore, we propose to approximate the Jacobian inverse in the Newton step with the Jacobian matrix itself.
+We show that this is a valid approximation that (a) is fundamentally different from the gradient descent direction used in previous works, (b) is cheap to compute, with only two forward passes through the denoiser required, and (c) can quickly converge to the desired solution in large pre-trained diffusion models.
+To evaluate, we generate images under both linear and non-linear constraints.
+We first show that our approach matches the results of state-of-the-art methods on free-form inpainting and $8\times$ super-resolution at a fraction of the inference time.
+We then demonstrate how existing methods fail at inpainting large regions, while our algorithm obtains results closer to a fully fine-tuned diffusion model on inpainting.
+Finally, we show how we can apply our algorithm to non-linear constraints and perform style-guided and mask-guided generation, where the proposed method consistently generates images that satisfy the constraints better than existing approaches.
+Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
+Constraint PSLD - 8min FreeDoM - 1min MPGD - 18s Ours - 15s Ground Truth
+[28] [44] [13]
+Figure 1: When tasked with completing the missing half of an image, previous methods are slow and fail to capture the important long-range dependencies between pixels.
+The proposed algorithm generates a reasonable image at a fraction of the time.
 2 Background
 2.1 Denoising Diffusion Models
-Denoising diffusion models, proposed for image generation by Ho et al.
-(2020), have now been widely adopted mainly due to their exceptional ability to synthesize diverse and high-quality samples.
-The original formulation views the training and inference process as a hierarchical latent variable model ${\bm{x}}_{T}\rightarrow{\bm{x}}_{T-1}\rightarrow\dots\rightarrow{\bm{x}}_{1}\rightarrow{\bm{x}}_{0}$, where the final latent is distributed normally ${\bm{x}}_{T}~{}\sim N(\mathbf{0},{\bm{I}})$ and $p({\bm{x}}_{0})$ represents the data distribution.
-After choosing a noise schedule $a_{t}$ that defines the forward transitions ${\bm{x}}_{t}\rightarrow{\bm{x}}_{t+1}$ that corrupt the data with Gaussian noise, usually centered at $\sqrt{\frac{a_{t+1}}{a_{t}}}{\bm{x}}_{t}$ and with variance $(1-\frac{a_{t+1}}{a_{t}})$, the model is trained to reverse each individual step in the diffusion process by predicting the noise added to the clean sample.
-In this formulation, the predicted noise is shown to approximate the score function $\nabla_{{\bm{x}}_{t}}\log p_{t}({\bm{x}}_{t})$ (Song et al., 2021) of the diffusion latent variables.
-Further iterations of denoising diffusion introduced classifier guidance (Dhariwal & Nichol, 2021), which adapted a pre-trained unconditional diffusion model for conditional sampling by training an additional classifier $p({\bm{y}}|{\bm{x}}_{t})$.
-During inference, each reverse step is also perturbed by the gradient $\nabla_{{\bm{x}}_{t}}\log p({\bm{y}}|{\bm{x}}_{t})$ which guides the diffusion latent ${\bm{x}}_{t}$ towards regions that also satisfy the condition ${\bm{y}}$.
-Classifier-free guidance (Ho & Salimans, 2022) eliminated the need for training an additional classifier to perform conditional generation by incorporating it into the base diffusion model training process.
-Latent diffusion models (LDMs) proposed Rombach et al.
-(2022), reduced the complexity of training large diffusion models by training the diffusion in the latent space of a pre-trained autoencoder.
-By utilizing an encoder $\mathcal{E}$ and decoder $\mathcal{D}$ that accurately reconstruct images ${\bm{x}}_{0}\approx\mathcal{D}(\mathcal{E}({\bm{x}}_{0}))$, the diffusion process can be made more efficient as the learned redundant information in an image is left for the decoder to reconstruct.
-Most large text-to-image diffusion models, such as Stable Diffusion (Podell et al., 2024), are based on this encoder-decoder approach to reduce complexity.
+Denoising diffusion models [32, 16] have been widely adopted due to their exceptional ability to synthesize diverse and high-quality samples.
+The original formulation treats the training and inference process as a hierarchical latent variable model ${\bm{x}}_{T}\rightarrow{\bm{x}}_{T-1}\rightarrow\dots\rightarrow{\bm{x}}_{1}\rightarrow{\bm{x}}_{0}$, where the final latent is distributed normally ${\bm{x}}_{T}~\sim N(\mathbf{0},{\bm{I}})$ and $p({\bm{x}}_{0})$ represents the data distribution.
+Given a noise schedule $\alpha_{t}$ that defines the forward transitions ${\bm{x}}_{t}\rightarrow{\bm{x}}_{t+1}$ that corrupt the data with Gaussian noise, usually centered at $\sqrt{\frac{\alpha_{t+1}}{\alpha_{t}}}{\bm{x}}_{t}$ and with variance $(1-\frac{\alpha_{t+1}}{\alpha_{t}})$, the model is trained to reverse each step in the diffusion process by predicting the noise added to the clean sample.
+The predicted noise is shown to approximate the score function $\nabla_{{\bm{x}}_{t}}\log p_{t}({\bm{x}}_{t})$ [34] of the diffusion latent variables.
+Further iterations of denoising diffusion introduced classifier guidance [9], which adapted a pre-trained unconditional diffusion model for conditional sampling by training an additional classifier $p({\bm{y}}|{\bm{x}}_{t})$.
+During inference, each reverse step also includes the gradient $\nabla_{{\bm{x}}_{t}}\log p({\bm{y}}|{\bm{x}}_{t})$, which guides the diffusion latent ${\bm{x}}_{t}$ towards regions that also satisfy the condition $\bm{y}$}.
+However, if the conditioning $\bm{y}$} is already given, training this additional classifier on top of the diffusion model is costly.
+Classifier-free guidance [15] eliminated the need for an additional classifier by incorporating the conditional guidance into the base diffusion model training process.
+The most widely adopted formulation of denoising diffusion is Latent diffusion models (LDMs) [27].
+LDMs reduce the complexity by modeling the compressed latent space of an autoencoder instead of images.
+With an encoder $\mathcal{E}$ and decoder $\mathcal{D}$ that accurately reconstruct images ${\bm{x}}_{0}\approx\mathcal{D}(\mathcal{E}({\bm{x}}_{0}))$, the diffusion process can be made more efficient as redundant information in an image is left for the decoder to reconstruct.
+Most large text-to-image diffusion models, such as Stable Diffusion [27, 23], are based on the LDM approach.
 2.2 Gradient descent steps for constrained sampling
-Previous works have studied whether large pre-trained diffusion models, which required significant investment to train, can be directly used for inference under novel conditions without additional tuning for each different constraint (Chung et al., 2023).
-The most typical problem formulation is denoising the sequence ${\bm{x}}_{T},{\bm{x}}_{T-1},...,{\bm{x}}_{1},{\bm{x}}_{0}$ under the linear constraint on the final signal in the form ${\bm{A}}{\bm{x}}_{0}={\bm{y}}$, or using a relaxed version that requires minimization of $||{\bm{A}}{\bm{x}}_{0}-{\bm{y}}||_{2}^{2}$, possibly as part of the likelihood function $p({\bm{y}}|{\bm{x}}_{0})={\cal N}({\bm{y}};{\bm{A}}{\bm{x}}_{0},\sigma^{2}{\bm{I}})$.
-In contrast to classifier guidance, which explicitly trained a separate model for the likelihood $p({\bm{y}}\mid{\bm{x}}_{t})$, the linear constraint only applies to the final, noise-free image ${\bm{x}}_{0}$.
-Thus, existing methods rely on Tweedie’s formula (Efron, 2011), by which denoising diffusion models approximating $\nabla_{{\bm{x}}_{t}}\log p_{t}({\bm{x}}_{t})$ can be used to express the expected value of ${\bm{x}}_{0}$, denoted as $\hat{{\bm{x}}}_{0}$.
-Including the constraint as if an additional observed variable ${\bm{y}}$ was generated requires the addition of the term $\nabla_{{\bm{x}}_{t}}\log p({\bm{y}}\mid{\bm{x}}_{t})$ to every diffusion sampling step, and most previous methods focused on different approximations of $p({\bm{y}}\mid{\bm{x}}_{t})$ using the estimated $\hat{{\bm{x}}}_{0}$.
+Previous works studied whether large pre-trained diffusion models, which required significant investment to train, can be directly used for inference under novel conditions without additional tuning for each different constraint [5].
+The typical problem formulation is denoising the sequence ${\bm{x}}_{T},{\bm{x}}_{T-1},...,{\bm{x}}_{1},{\bm{x}}_{0}$ under a linear constraint on the final signal in the form ${\bm{A}}{\bm{x}}_{0}={\bm{y}}$, or a relaxed version that minimizes ${\bm{A}}{\bm{x}}_{0}-{\bm{y}}||_{2}^{2}$ as part of the likelihood $p({\bm{y}}|{\bm{x}}_{0})={\cal N}({\bm{y}};{\bm{A}}{\bm{x}}_{0},\sigma^{2}{\bm{I}})$.
+Contrary to classifier guidance, which trained a separate model for the likelihood $p({\bm{y}}\mid{\bm{x}}_{t})$, the linear constraint only applies to the final, noise-free image ${\bm{x}}_{0}$.
+Thus, existing methods rely on Tweedie’s formula [10], by which denoising diffusion models approximating $\nabla_{{\bm{x}}_{t}}\log p_{t}({\bm{x}}_{t})$ can be used to express the expected value of ${\bm{x}}_{0}$, denoted as $\hat{{\bm{x}}}_{0}$.
+Including the constraint as if an additional observed variable $\bm{y}$} was generated requires adding $\nabla_{{\bm{x}}_{t}}\log p({\bm{y}}\mid{\bm{x}}_{t})$ to every diffusion sampling step, and previous works considered different approximations of $p({\bm{y}}\mid{\bm{x}}_{t})$ using the estimated $\hat{{\bm{x}}}_{0}$ [5, 44].
 Regardless of how the constraint gradient is applied, the regular denoising diffusion steps are altered so that at each $t$, the generated latent ${\bm{x}}_{t}$ is moved in the direction reducing the cost
 $C({\bm{x}}_{t})=({\bm{A}}\hat{{\bm{x}}}_{0}({\bm{x}}_{t})-{\bm{y}})^{T}({\bm{A}}\hat{{\bm{x}}}_{0}({\bm{x}}_{t})-{\bm{y}}).$ (1)
-For example, in the case of inpainting missing pixels, the matrix ${\bm{A}}$ extracts the subsection of the known pixels in image ${\bm{x}}_{0}$ to be compared with a given target ${\bm{y}}$.
+For example, in the case of inpainting missing pixels, the matrix $\bm{A}$} extracts the subsection of the known pixels in image $\hat{{\bm{x}}}_{0}$ to be compared with a given target $\bm{y}$}.
 The estimated expected value of ${\bm{x}}_{0}$ at the end of the chain is provided by the diffusion model as a nonlinear function $\hat{{\bm{x}}}_{0}({\bm{x}}_{t})$ learned by the denoiser network during training.
 Typically, these moves are gradient descent moves, i.e.
-moves of ${\bm{x}}_{t}$ in the direction of $-{\bm{e}}_{t}$ where
-$-{\bm{e}}_{t}=-\nabla_{{\bm{x}}_{t}}C({\bm{x}}_{t})=-{\bm{J}}^{T}{\bm{A}}^{T}({\bm{A}}\hat{{\bm{x}}}_{0}-{\bm{y}})=-{\bm{J}}^{T}{\bm{e}},$ (2)
-${\bm{J}}=\nabla_{{\bm{x}}_{t}}\hat{{\bm{x}}}_{0}({\bm{x}}_{t}),\quad{\bm{e}}={\bm{A}}^{T}({\bm{A}}\hat{{\bm{x}}}_{0}-{\bm{y}}).$ (3)
-Note that we name the error signal ${\bm{e}}$ as the (negative) direction of the gradient w.r.t. ${\bm{x}}_{0}$ itself, and the ${\bm{e}}_{t}$ is the matching move in ${\bm{x}}_{t}$.
+moves of ${\bm{x}}_{t}$ in the direction of ${\bm{e}}_{t}$ where
+${\bm{e}}_{t}=\nabla_{{\bm{x}}_{t}}C({\bm{x}}_{t})={\bm{J}}^{T}{\bm{A}}^{T}({\bm{A}}\hat{{\bm{x}}}_{0}-{\bm{y}})={\bm{J}}^{T}{\bm{e}},$ (2)
+${\bm{J}}=\nabla_{{\bm{x}}_{t}}\hat{{\bm{x}}}_{0}({\bm{x}}_{t}),\quad{\bm{e}}={\bm{A}}^{T}({\bm{A}}\hat{{\bm{x}}}_{0}-{\bm{y}}).$
+Note that we name the error signal $\bm{e}$} as the (negative) direction of the gradient w.r.t. ${\bm{x}}_{0}$ itself, and the ${\bm{e}}_{t}$ is the matching move in the noisy ${\bm{x}}_{t}$.
 The matrix ${\bm{A}}^{T}$ inverses the operation of $\bm{A}$} and can usually be computed for each task a-priori.
-However, computation of the full Jacobian ${\bm{J}}$ is impractical, so the gradient is instead computed using backpropagation through $C({\bm{x}}_{t})$.
-In (Chung et al., 2023) and (Chung et al., 2024) for example, gradient steps of this form are applied to ${\bm{x}}_{t}$ at each step $t$ of generation before moving on to the next stage.
-As optimization of ${\bm{x}}_{t}$ might reduce the total noise in the image below what the denoising at $t-1$ was trained for, the gradient steps moving ${\bm{x}}_{t}$ towards optimizing $C({\bm{x}}_{t})$ can be combined with adding additional noise, which could also be seen as a form of stochastic averaging, as done by Yu et al.
-(2023).
-3 Method
-3.1 Approximate Newton steps
+Since computing the full Jacobian $\bm{J}$} is impractical, the gradient is instead computed using backpropagation through $C({\bm{x}}_{t})$, which yields the direction ${\bm{J}}^{T}{\bm{e}}$.
+Chung et al.
+[5], [6], for example, apply gradient steps of this form to ${\bm{x}}_{t}$ at each step $t$ of generation before moving on to the next stage.
+As optimization of ${\bm{x}}_{t}$ might reduce the total noise in the image below what the denoising at t1t-1 was trained for, the gradient steps moving ${\bm{x}}_{t}$ towards optimizing $C({\bm{x}}_{t})$ can be combined with adding additional noise, which could also be seen as a form of stochastic averaging, as done by Yu et al.
+[44].
+3 Method: Approximate Newton steps
 We start by observing that a Newton optimization step, instead of moving ${\bm{x}}_{t}$ in the gradient descent direction ${\bm{J}}^{T}{\bm{e}}$, moves it in the direction ${\bm{J}}^{-1}{\bm{e}}$.
-We demonstrate that on a more general form of the cost
+We demonstrate it on a more general form of the cost
 $C({\bm{x}}_{t})=(f(\hat{{\bm{x}}}_{0}({\bm{x}}_{t}))-{\bm{y}})^{T}(f(\hat{{\bm{x}}}_{0}({\bm{x}}_{t}))-{\bm{y}}),$ (3)
-for some target ${\bm{y}}$ to be matched with projection function $f$, which in the linear case is $f({\bm{x}})={\bm{A}}{\bm{x}}$. The Gauss-Newton optimization would first approximate
+for some target $\bm{y}$} to be matched with projection function $f$, which in the linear case is $f({\bm{x}})={\bm{A}}{\bm{x}}$. The Newton optimization would first approximate
 $f(\hat{{\bm{x}}}_{0}({\bm{x}}_{t}-{\bm{e}}_{t}))\approx f(\hat{{\bm{x}}}_{0}({\bm{x}}_{t}))-{\bm{J}}_{f}{\bm{J}}{\bm{e}}_{t},$ (4)
-where ${\bm{J}}_{f}$ is the Jacobian of $f$. We then rewrite the cost of move $-{\bm{e}}_{t}$ as
-$C({\bm{x}}_{t}-{\bm{e}}_{t})=$ $(f(\hat{{\bm{x}}}_{0}({\bm{x}}_{t}))-{\bm{J}}_{f}{\bm{J}}{\bm{e}}_{t}-{\bm{y}})^{T}$
-$(f(\hat{{\bm{x}}}_{0}({\bm{x}}_{t}))-{\bm{J}}_{f}{\bm{J}}{\bm{e}}_{t}-{\bm{y}}).$ (6)
+where ${\bm{J}}_{f}$ is the Jacobian of $f$. We then rewrite the cost of move ${\bm{e}}_{t}$ as
+$C({\bm{x}}_{t}-{\bm{e}}_{t})=(f(\hat{{\bm{x}}}_{0}({\bm{x}}_{t}))-{\bm{J}}_{f}{\bm{J}}{\bm{e}}_{t}-{\bm{y}})^{T}(f(\hat{{\bm{x}}}_{0}({\bm{x}}_{t}))-{\bm{J}}_{f}{\bm{J}}{\bm{e}}_{t}-{\bm{y}}).$ (5)
 The cost is minimized by setting $\nabla_{{\bm{e}}_{t}}C=0$ to get the system
-${\bm{J}}^{T}{\bm{J}}_{f}^{T}(f(\hat{{\bm{x}}}_{0}({\bm{x}}_{t}))-{\bm{y}})={\bm{J}}^{T}{\bm{J}}_{f}^{T}{\bm{J}}_{f}{\bm{J}}{\bm{e}}_{t}$
-$\Leftrightarrow{\bm{e}}_{t}={\bm{J}}^{-1}{\bm{J}}_{f}^{-1}(f(\hat{{\bm{x}}}_{0}({\bm{x}}_{t}))-{\bm{y}})$ (7)
+${\bm{J}}^{T}{\bm{J}}_{f}^{T}(f(\hat{{\bm{x}}}_{0}({\bm{x}}_{t}))-{\bm{y}})={\bm{J}}^{T}{\bm{J}}_{f}^{T}{\bm{J}}_{f}{\bm{J}}{\bm{e}}_{t}\Leftrightarrow{\bm{e}}_{t}={\bm{J}}^{-1}{\bm{J}}_{f}^{-1}(f(\hat{{\bm{x}}}_{0}({\bm{x}}_{t}))-{\bm{y}})$ (6)
 ${\bm{e}}_{t}={\bm{J}}^{-1}{\bm{e}},\quad{\bm{e}}={\bm{J}}_{f}^{-1}(f(\hat{{\bm{x}}}_{0}({\bm{x}}_{t}))-{\bm{y}}),$ (7)
 assuming the inverses exist.
-In the case of $f({\bm{x}})={\bm{A}}{\bm{x}}$ for inpainting and superresolution tasks, where ${\bm{y}}$ is lower-dimensional, the inverse of ${\bm{J}}_{f}$, obviously is not defined, but we can use the pseudoinverse.
+In the case of $f({\bm{x}})={\bm{A}}{\bm{x}}$ for inpainting and super-resolution tasks, where $\bm{y}$} is lower-dimensional, the inverse of ${\bm{J}}_{f}$, is not defined, but we can use the pseudo-inverse.
 Similarly, for a non-linear $f$, we can compute $\bm{e}$} using numerical methods, e.g.
-backpropagation through $f$ when it is a neural network.
-Therefore, ${\bm{e}}$ is the same in both gradient descent and Newton optimization, but the relationship to ${\bm{e}}_{t}$ is different:
-$\text{GD}:\ {\bm{e}}_{t}={\bm{J}}^{T}{\bm{e}},\ \ \text{Newton}:\ {\bm{J}}{\bm{e}}_{t}={\bm{e}}$ (9)
-The gradient descent step can be computed without directly evaluating the Jacobian by back-propagation on the scalar cost $C$.
-For the computation of ${\bm{J}}^{-1}{\bm{e}}$ on the other hand, we have no such method.
-However, we can efficiently compute the move
-$\epsilon{\bm{e}}_{t}=\epsilon{\bm{J}}{\bm{e}}.$ (10)
-We follow the Newton recipe suggesting multiplication of the error by ${\bm{J}}$ in Eq.
-(9) but multiply ${\bm{e}}$ by it instead of ${\bm{e}}_{t}$.
-Intuitively, this is justified by the fact that the mode of ${\bm{x}}_{t-1}$ in the next stage of denoising is a weighted sum of $\hat{{\bm{x}}_{0}}$ and ${\bm{x}}_{t}$, meaning that the changes in ${\bm{x}}_{t}$ can be equivalently represented by proportional moves in the same direction of either component.
-This is the consequence of the modeling assumption that in the posterior the stages are separated by independent Gaussian noise, and early inpaiting models simply assumed ${\bm{e}}_{t}={\bm{e}}$, missing on the effect of the image prior $p({\bm{x}}_{0})$, which is implicit in the learned function ${\bm{x}}_{0}({\bm{x}}_{t})$.
-In addition to that, in Appendix A.1 we provide an analysis of the convergence on the algorithm using inexact Newton methods (Dembo et al., 1982).
+backpropagation through $f$ when it is a neural network, which would approximate $\bm{e}$} using ${\bm{J}}_{f}^{T}$.
+Therefore, $\bm{e}$} is the same in gradient descent and Newton optimization, but its relationship to ${\bm{e}}_{t}$ differs
+$\text{(a)}\ \text{GD}:\ {\bm{e}}_{t}={\bm{J}}^{T}{\bm{e}},\quad\text{(b)}\ \text{Newton}:\ {\bm{J}}{\bm{e}}_{t}={\bm{e}}.$ (8)
+Gradient descent can be computed without directly evaluating the Jacobian by backpropagation on the scalar cost $C$.
+For the computation of ${\bm{J}}^{-1}{\bm{e}}$, on the other hand, we have no such method.
+In cases where computing the inverse of the Jacobian is prohibitive, inexact Newton methods [7] propose first finding an approximation ${\bm{e}}_{t}^{*}$ to the solution of Eq (8) (b), performing the Newton step using the approximate solution, and reiterating until convergence.
+When the residual $\bm{r}$} is strictly reduced at every step, inexact Newton methods converge to the correct solution
+${\bm{r}}={\bm{e}}-{\bm{J}}{\bm{e}}_{t}^{*},\quad\frac{\lVert{\bm{r}}\rVert_{2}}{\lVert{\bm{e}}\rVert_{2}}<\eta,\ \eta\in[0,1).$ (9)
+Since we know that the input and output of the denoiser are related by additive Gaussian noise, we propose the inexact move
+${\bm{e}}_{t}={\bm{J}}{\bm{e}}$ (10)
+where instead of computing the Jacobian inverse, we use the Jacobian-error vector product.
+If the Jacobian tells us how to transform a local change in the input to a change in the output, then we posit that, for denoising diffusion models, we can use the same transformation for the inverse.
+Substituting ${\bm{e}}_{t}^{*}$ with the proposed update, adding a learning rate $\lambda$ to control convergence, we get the residual
+${\bm{r}}={\bm{e}}-\lambda{\bm{J}}^{2}{\bm{e}}=({\bm{I}}-\lambda{\bm{J}}^{2}){\bm{e}}.$ (11)
+For this residual to be strictly reduced at every iteration of our algorithm, we require
+$\frac{\lVert{\bm{r}}\rVert_{2}}{\lVert{\bm{e}}\rVert_{2}}=\frac{\lVert({\bm{I}}-\lambda{\bm{J}}^{2}){\bm{e}}\rVert_{2}}{\lVert{\bm{e}}\rVert_{2}}\leq\frac{\lVert{\bm{I}}-\lambda{\bm{J}}^{2}\rVert_{2}\lVert{\bm{e}}\rVert_{2}}{\lVert{\bm{e}}\rVert_{2}}=\lVert{\bm{I}}-\lambda{\bm{J}}^{2}\rVert_{2}<\eta.$ (12)
+As long as the spectrum of the denoiser Jacobian matrix is bounded, we can choose a small enough $\lambda$ that strictly reduces the residual.
+However, the proposed move would converge slowly if the only $\lambda$ allowed were very small.
+In practice, we find that we can use large learning rates, making the convergence similar or even better than gradient descent.
+In Appendix A.1, we present an analysis of the spectral properties of the Jacobian matrix of the Stable Diffusion denoiser used in our experiments.
+Furthermore, in Appendix A.3 we demonstrate the differences between our proposed inexact and the exact Newton method on a small-scale experiment.
+Figure 2: We showcase the difference between the proposed method to compute the update direction ($\bm{J}$}) and gradient descent (${\bm{J}}^{T}$).
+The heatmaps indicate where the input ${\bm{x}}_{t}$ would change when perturbing a single pixel in the output, denoted in green.
+The two directions are considerably different, with ours capturing better longer-range correlations and maintaining shapes.
+Even though we use finite differences, the direction computed from $\bm{J}$} is sharper in some regions, like the outlines.
 The direction of optimization we propose in Eq.
-(10) has at least two advantages over the usual gradient descent update.
-First, the direction $\epsilon{\bm{e}}_{t}$ can be computed numerically to save both on computation (2-fold) and memory (2.5-fold) compared to using backpropagation on the cost in Eq.
+(10) has another advantage over the gradient descent update.
+The direction ${\bm{e}}_{t}$ can be computed numerically to save both on computation and memory compared to using backpropagation on the cost in Eq.
 (1).
-To derive that update, consider the function $f(s)=\hat{{\bm{x}}}_{0}({\bm{x}}_{t}-s{\bm{e}})$ where the variable $s$ is scalar.
+To derive the update, consider the function $h(s)=\hat{{\bm{x}}}_{0}({\bm{x}}_{t}+s{\bm{e}})$ where the variable $s$ is scalar.
 Its derivative at $s=0$ is
-$\frac{df}{ds}=-{\bm{J}}{\bm{e}},$ (11)
-and so the direction ${\bm{h}}$ can be approximated numerically using the approximate derivative of $f(s)$ at $s=0$
-$f^{\prime}=\frac{1}{\delta}[\hat{{\bm{x}}}_{0}({\bm{x}}_{t}+\delta{\bm{e}})-\hat{{\bm{x}}}_{0}({\bm{x}}_{t})],$ (12)
-${\bm{e}}_{t}=-\epsilon{\bm{J}}{\bm{e}}\approx\frac{\epsilon}{\delta}[\hat{{\bm{x}}}_{0}({\bm{x}}_{t}+\delta{\bm{e}})-\hat{{\bm{x}}}_{0}({\bm{x}}_{t})],$ (13)
-requiring no backpropagation but instead two forward passes through the network: one to compute $\hat{{\bm{x}}}_{0}({\bm{x}}_{t})$ and another to compute $\hat{{\bm{x}}}_{0}$ for the ${\bm{x}}_{t}$ perturbed in the negative direction of the error vector.
-The second advantage comes from the asymmetry of the Jacobian ${\bm{J}}$.
-The theoretically optimal ${\bm{x}}_{0}({\bm{x}}_{t})$ would have a symmetric Jacobian, as $\hat{{\bm{x}}}_{0}$ approximates the true expectation $E[{\bm{x}}_{0}|{\bm{x}}_{t}]=\frac{1}{\sqrt{\alpha_{t}}}[{\bm{x}}_{t}+{\bm{v}}_{t}\nabla_{{\bm{x}}_{t}}\log p_{t}({\bm{x}}_{t})]$, and the gradient of this is indeed symmetric:
+$\left.\frac{dh}{ds}\right\rvert_{s=0}={\bm{J}}{\bm{e}},$ (13)
+and so the direction ${\bm{e}}_{t}$ can be computed using the numerical derivative of $h(s)$ at $s=0$
+${\bm{e}}_{t}=\lambda{\bm{J}}{\bm{e}}=\lambda\left.\frac{dh}{ds}\right\rvert_{s=0}\approx\lambda\frac{h(\delta)-h(0)}{\delta}=\lambda\frac{\hat{{\bm{x}}}_{0}({\bm{x}}_{t}+\delta{\bm{e}})-\hat{{\bm{x}}}_{0}({\bm{x}}_{t})}{\delta},$ (14)
+requiring no backpropagation but instead two forward passes through the network: one to compute $\hat{{\bm{x}}}_{0}({\bm{x}}_{t})$ and another to compute $\hat{{\bm{x}}}_{0}$ for the ${\bm{x}}_{t}$ perturbed in the direction of the error vector.
+It is important to point out that the theoretically optimal ${\bm{x}}_{0}({\bm{x}}_{t})$ would have a symmetric Jacobian, as $\hat{{\bm{x}}}_{0}$ approximates the true expectation $E[{\bm{x}}_{0}|{\bm{x}}_{t}]=\frac{1}{\sqrt{\alpha_{t}}}[{\bm{x}}_{t}+{\bm{v}}_{t}\nabla_{{\bm{x}}_{t}}\log p_{t}({\bm{x}}_{t})]$, and the gradient of this is indeed symmetric
 $\nabla_{{\bm{x}}_{t}}E[{\bm{x}}_{0}|{\bm{x}}_{t}]=\frac{1}{\sqrt{\alpha_{t}}}[{\bm{I}}+{\bm{v}}_{t}\nabla^{2}\log p_{t}({\bm{x}}_{t})].$ (15)
 This would render the gradient descent and the proposed update equivalent.
-In Section 3.2, we show that in practice, the trained denoiser models do not satisfy this ideal condition, making the two update directions noticeably different.
-Whether the denoising diffusion models are trained with score matching in mind (Song et al., 2021) or using the variational method of Ho et al.
-(2020), they do not directly optimize to match the real score $\nabla_{{\bm{x}}_{t}}\log p_{t}({\bm{x}}_{t})$ everywhere nor are they constrained to produce symmetric Jacobians.
-We hypothesize that this discrepancy in the two update directions is one of the reasons why our optimization of ${\bm{x}}_{t}$ may be more suitable for some applications, as demonstrated in Section 4.
+Yet, we find that the trained denoiser models do not satisfy this ideal condition, making the two update directions noticeably different.
+In Figure 2, we visualize this difference between $\bm{J}$} and ${\bm{J}}^{T}$ by computing rows of $\bm{J}$} and ${\bm{J}}^{T}$, and averaging the magnitude of all matrix values that contribute to each pixel.
+This indicates that there is a difference in which pixels in ${\bm{x}}_{t}$ would change when perturbing a single pixel in $\hat{{\bm{x}}}_{0}({\bm{x}}_{t})$ for the two update directions.
+Further experiments on the difference between the two update directions are included in Appendix A.2.
+Whether the denoising diffusion models are trained with score matching in mind [34] or using the variational method of Ho et al.
+[16], they do not directly optimize to match the real score $\nabla_{{\bm{x}}_{t}}\log p_{t}({\bm{x}}_{t})$ everywhere nor are they constrained to produce symmetric Jacobians.
+We hypothesize that this discrepancy in the two update directions is another reason why our optimization of ${\bm{x}}_{t}$ may be more suitable for some applications, which we demonstrate in the experiments (Section 4).
 Lastly, the proposed update of Eq.
-(10) only requires us to provide the direction of the error ${\bm{e}}$, which points locally towards a lower cost for our constraint on ${\bm{x}}_{0}$.
-As already discussed, for linear constraints such as inpainting, this direction can be obtained in closed form using the inverse of the operator expressed by ${\bm{A}}$.
-For non-linear constraints, such as style-guidance, ${\bm{e}}$ can be computed using backpropagation through the differentiable non-linear constraint model.
-Note that this backpropagation operation does not backpropagate through the denoiser itself, but only through the constraint w.r.t.
-the estimated $\hat{{\bm{x}}}_{0}$.
-A general version of the proposed algorithm for linear and non-linear constraints is presented in Algorithm 1.
-We choose DDIM as our sampling algorithm (Song et al., 2020).
-Applying other sampling methods should be intuitive by interleaving the diffusion and our gradient updates on ${\bm{x}}_{t}$.
-Algorithm 1 The proposed algorithm for constrained sampling using a pre-trained diffusion model.
-1:  Input: Pre-trained diffusion model $\hat{{\bm{x}}}_{0}({\bm{x}}_{t})$, constraint $C({\bm{x}}_{0},{\bm{y}})=\lVert f(\hat{{\bm{x}}}_{0}({\bm{x}}_{t})-y\rVert_{2}^{2}$, condition ${\bm{y}}$, gradient step size $\delta$, iterations $K$, learning rate $\lambda$, diffusion step size $s$ and schedule parameters $\alpha_{i},\beta_{i}$
-2:  ${\bm{x}}_{T}\sim N(\textbf{0},{\bm{I}})$
-3:  for $t\in\{T,T-s,T-2s,\dots,s\}$ do
-4:     for $i\in\{1,2,\dots,K\}$ do
-5:        ${\bm{e}}={\bm{J}}_{f}^{-1}(f(\hat{{\bm{x}}}_{0}({\bm{x}}_{t}))-{\bm{y}})$Task-dependent}
-6:        ${\bm{e}}_{t}=[\hat{{\bm{x}}}_{0}({\bm{x}}_{t}+\delta{\bm{e}})-\hat{{\bm{x}}}_{0}({\bm{x}}_{t})]/\delta$
-7:        ${\bm{x}}_{t}={\bm{x}}_{t}+\lambda{\bm{e}}_{t}$
-8:     end for
-9:     ${\bm{z}}_{t}\sim N(\textbf{0},{\bm{I}})$
-10:     $\mathbf{\epsilon}_{t}=\frac{1}{\sqrt{1-\alpha_{t-s}}}{\bm{x}}_{t}-\frac{\sqrt{\alpha_{t-s}}}{\sqrt{1-\alpha_{t-s}}}\hat{{\bm{x}}}_{0}({\bm{x}}_{t})$
-11:     ${\bm{x}}_{t-s}=\sqrt{\alpha_{t-s}}\hat{{\bm{x}}}_{0}({\bm{x}}_{t})+\sqrt{1-\alpha_{t-s}}\mathbf{\epsilon}_{t}+\beta_{t-s}{\bm{z}}_{t}$
+(10) only requires us to provide the direction of the error $\bm{e}$}, which points locally towards a lower cost for the constraint on ${\bm{x}}_{0}$.
+For linear constraints such as inpainting, this direction can be obtained in closed form using the inverse of the operator expressed by $\bm{A}$}.
+For non-linear constraints, $\bm{e}$} does not have to be computed in closed form from ${\bm{J}}_{f}^{-1}$, and can be approximated.
+In our experiments, we use backpropagation through the differentiable non-linear VAE decoder and constraint, which is cheaper than backpropagating through the denoiser.
+An alternative would be to utilize another inexact Newton approximation to avoid backpropagating through the constraint altogether.
+We discuss this in Appendix A.4, where for linear constraints applied to the the VAE output space (i.e.
+pixels), we propose an inexact Newton method that utilizes the Jacobian of the VAE encoder to approximate the inverse of the VAE decoder Jacobian.
+Similar approximations have been discussed in the literature before [36] and can be particularly useful when backpropagation is infeasible [41].
+The proposed method for sampling with linear and non-linear constraints is described in Algorithm 1, using DDIM as the diffusion sampling algorithm [33].
+Using other diffusion sampling algorithms [19] is intuitive by interleaving the diffusion and our constraint gradient updates on ${\bm{x}}_{t}$ (Appendix A.6).
+Algorithm 1 The proposed algorithm for sampling under linear and non-linear constraints.
+1:Input: Pre-trained diffusion model $\hat{{\bm{x}}}_{0}({\bm{x}}_{t})$, linear constraint $C({\bm{x}}_{0},{\bm{y}})=\lVert{\bm{A}}{{\bm{x}}}_{0}({\bm{x}}_{t})-{\bm{y}}\rVert_{2}^{2}$ or non-linear constraint $C({\bm{x}}_{0},{\bm{y}})=\lVert f({{\bm{x}}}_{0}({\bm{x}}_{t}))-{\bm{y}}\rVert_{2}^{2}$, condition $\bm{y}$}, step size $\delta$, iterations $K$, learning rate $\lambda$, diffusion step size $s$ and schedule parameters $\alpha_{i},\beta_{i}$
+[…]
+3:for $t=T,T-s,T-2s,\dots,s$ do
+4:  for $i=1,2,\dots,K$ do
+5:   if Linear then
+6:    ${\bm{e}}={\bm{A}}^{T}({\bm{A}}\hat{{\bm{x}}}_{0}({\bm{x}}_{t})-{\bm{y}})$ {Using pseudoinverse ${\bm{A}}^{T}$}
+7:   else if Non-linear then
+8:    ${\bm{e}}={\bm{J}}_{f}^{T}(f(\hat{{\bm{x}}}_{0}({\bm{x}}_{t}))-{\bm{y}})=\nabla_{\hat{{\bm{x}}}_{0}}C(\hat{{\bm{x}}}_{0},{\bm{y}})$ {Using backpropagation through ff}
+9:   end if
+10:   ${\bm{e}}_{t}=[\hat{{\bm{x}}}_{0}({\bm{x}}_{t}+\delta{\bm{e}})-\hat{{\bm{x}}}_{0}({\bm{x}}_{t})]/\delta$
+11:   ${\bm{x}}_{t}={\bm{x}}_{t}-\lambda{\bm{e}}_{t}$
 12:  end for
-13:  Return: ${\bm{x}}_{0}$
-3.2 Symmetry of the Jacobian and difference in updates
-Figure 2: Sample pairs $(i,j),(k,l)$ of the denoiser Jacobian $\nabla_{{\bm{x}}_{t}}\hat{{\bm{x}}_{0}}=\nabla_{{\bm{x}}_{t}}E[{\bm{x}}_{0}|{\bm{x}}_{t}]$ for different timesteps $t$.
-We observe that the Jacobian is not symmetric, justifying the difference between the proposed update steps and the previously used gradient descent steps.
-Although, theoretically, the Jacobian of the denoiser should be symmetric (Eq.
-14), the trained diffusion model does not exactly match the real score and can yield non-symmetric Jacobians.
-In Figure 2, we verify this claim with a simple experiment; we select a random image from the ImageNet (Deng et al., 2009) validation set and employ the Stable Diffusion 1.5 model (Podell et al., 2024) to denoise at three different noise levels $t=\{900,700,400\}$.
-We encode the image using the VAE encoder, scale it and add appropriate noise to get the intermediate diffusion latent for each timestep.
-We then give the noisy image to the denoiser network and compute the gradients $\partial\hat{{\bm{x}}}_{0}^{k,l}/\partial{\bm{x}}_{t}^{i,j}$ and $\partial\hat{{\bm{x}}}_{0}^{i,j}/\partial{\bm{x}}_{t}^{k,l}$ for randomly chosen pixels $(i,j),(k,l)$ using backpropagation.
-When we plot the gradients and compute the correlation coefficient $r$ we observe that the values deviate from $y=x$, which would indicate a symmetric Jacobian.
-Figure 3: We present a simple experiment to compare the proposed gradient update ${\bm{J}}{\bm{e}}$ with that of previous methods ${\bm{J}}^{T}{\bm{e}}$.
-The goal is to update a noisy version of the image shown in (a) at $t=800$, such that the diffusion model’s prediction of the final image includes a red square in the middle, shown in (b).
-The results of gradient descent in (c) and our gradient update in (d) demonstrate that updates performed by previous methods lead to a considerably different result than the update we propose in this paper.
-Having shown that the Jacobian is not symmetric, we expect to find differences between the gradient updates of gradient descent (Eq.
-3) and our proposed update (Eq.
-10).
-To highlight these differences, we set up a toy experiment where we update an identical initial image with the two different directions, under the same condition.
-The experiment is showcased in Figure 3 and again utilizes the Stable Diffusion 1.5 model.
-We create a synthetic black-and-white grid image (Figure 3 (a)), which we blur, deterministically encode, scale and noise to get a diffusion latent at $t=800$.
-We then set up a constraint where we add a red square to the center of the original image (Figure 3 (b)).
-The brown region in the constraint indicates masked-out values, or equivalently, where we will not be applying a constraint.
-Instead of decoding the image and applying the constraint in image space, we also encode the constraint image and apply it directly in the Stable Diffusion latent space.
-We run 5 gradient updates with the same learning rate of 1 for each and demonstrate the final predicted $\hat{{\bm{x}}}_{0}$ for the ${\bm{J}}^{T}{\bm{e}}$ update of (3) (Figure 3 (c)) and the proposed ${\bm{J}}{\bm{e}}$ (10) (Figure 3 (d)).
-The final result shows how the model intends to change the entire image when asked to add a red square in the middle.
-We find that the resulting images differ substantially, with the proposed direction producing a more coherent image that tries to copy the newly introduced texture to the correct locations, i.e.
-the intersections of the lines.
-Although this is an empirical observation, we hypothesize that the two different directions can have vastly different effects on the image.
-In the Appendix A.2, we further visualize differences between ${\bm{J}}^{T}{\bm{e}}$ and ${\bm{J}}{\bm{e}}$, and in C.5, we repeat the previous experiment over multiple seeds.
+13:  ${\bm{z}}_{t}\sim N(\textbf{0},{\bm{I}})$
+14:  $\bm{\epsilon}_{t}=\frac{1}{\sqrt{1-\alpha_{t-s}}}{\bm{x}}_{t}-\frac{\sqrt{\alpha_{t-s}}}{\sqrt{1-\alpha_{t-s}}}\hat{{\bm{x}}}_{0}({\bm{x}}_{t})$
+15:  ${\bm{x}}_{t-s}=\sqrt{\alpha_{t-s}}\hat{{\bm{x}}}_{0}({\bm{x}}_{t})+\sqrt{1-\alpha_{t-s}-\beta^{2}_{t-s}}\bm{\epsilon}_{t}+\beta_{t-s}{\bm{z}}_{t}$ {DDIM step}
+16:end for
+17:Return: ${\bm{x}}_{0}$
 4 Experiments
 4.1 Linear Constraints
-Figure 4: Comparison between our method and existing algorithms on free-form inpainting and 8×8$\times$8 × super-resolution.
-We directly use the images and results from (Chung et al., 2024) since there is no code available to replicate their method.
-Inpaint (Free-form) SR (×$\times 8$)
-Method PSNR ↑ LPIPS ↓ FID ↓ PSNR ↑ LPIPS ↓ FID ↓ Time
-P2L (Chung et al., 2024) 21.99 0.229 32.82 23.38 0.386 51.81 30min*
-LDPS 21.54 0.332 46.72 23.21 0.475 61.09 6min
-PSLD (Rout et al., 2023) 20.92 0.251 40.57 23.17 0.471 60.81 8min
-Ours 21.73 0.258 19.39 24.26 0.455 60.99 15s/1min
-Table 1: Quantitative evaluation (PSNR, LPIPS, FID) on free-form inpainting and 8×8$\times$8 × superresolution on ImageNet.
-Our algorithm is competitive with the results of P2L which requires at least 30×30$\times$30 × more time for a single sample.
-Figure 5: Qualitative evaluation of large area (box) inpainting on ImageNet.
-Our method provides results close to the fine-tuned inpainting model while requiring a fraction of the time to run per image.
-We first verify our algorithm by generating images under linear constraints, which has been the main application of many previous algorithms (Chung et al., 2023; Rout et al., 2023; Chung et al., 2024).
+We first verify our algorithm by generating images under linear constraints, which has been the main application of many previous algorithms [5, 28, 6].
 We follow the evaluation setting of Saharia et al.
-(2022a) and test our method on ImageNet (Deng et al., 2009), using the first 1000 images from the 10k validation set of the ctest10k split.
-As our evaluation metrics, we measure the PSNR and LPIPS (Zhang et al., 2018) between the generated and real images, as well as the FID (Heusel et al., 2017) between the real and generated sets.
-We use Stable Diffusion 1.4, which is pre-trained on the LAION (Schuhmann et al., 2022) text-image pair dataset.
-All experiments were run on NVIDIA RTX A5000 24GB GPUs.
-Free-form Inpainting & Super-resolution
-We follow the previously described experiment settings of free-form inpainting and 8×8$\times$8 × super-resolution.
-For free-form inpainting, the masks are randomly sampled and mask out 10-20% of the image pixels, as done by Saharia et al.
-(2022a).
+[29] and test our method on ImageNet [8], using the first 1000 images from the 10k validation set of the ctest10k split.
+For evaluation, we measure the PSNR, LPIPS [46], and FID [14] between the real and generated images.
+We use Stable Diffusion 1.4, which is pre-trained on the LAION [31] text-image pair dataset.
+Experiments were run on an NVIDIA RTX A5000 24GB GPU.
+Free-form inpainting and $8\times$ super-resolution
+In free-form inpainting, masks are randomly sampled and mask out 10-20% of the image pixels.
 For inpainting with our method, we opt to directly operate on the Stable Diffusion latent given that Stable Diffusion VAE mostly compresses information locally.
@@
 We also apply a $3\times 3$ dilation kernel on the pixel mask before downsampling, masking out some extra pixels along the edge that we find the VAE fails to encode.
+For super-resolution, we cannot apply the constraint in the VAE latent space since image downsampling does not correspond to downsampling VAE latents.
+This makes super-resolution non-linear.
+It involves the differentiable decoder $\mathcal{D}$, and to compute the error direction $\bm{e}$}, we backpropagate the pixel-level linear constraint $({\bm{A}}\mathcal{D}(\hat{{\bm{x}}}_{0}))-{\bm{y}})^{T}({\bm{A}}\mathcal{D}(\hat{{\bm{x}}}_{0})-{\bm{y}})$ through the decoder network.
+We discuss backpropagation for computing the error direction further in the non-linear constraint experiments (4.2) but leave super-resolution in the linear section as done by previous works.
 For inpainting, we set the number of optimization steps $K=5$ over which we linearly decrease the learning rate $\lambda$ from 0.5 to 0.1.
 For super-resolution, we use $K=10$ and a constant $\lambda=0.1$.
-We perform only 20 DDIM (Song et al., 2020) steps for all experiments.
-Regarding super-resolution, we cannot apply the constraint in the VAE latent space as image pixel downsampling does not correspond to downsampling VAE latents.
-To compute the error direction ${\bm{e}}$ we backpropagate the pixel-level constraint cost $({\bm{A}}\mathcal{D}({\bm{x}}_{t})-{\bm{y}})^{T}({\bm{A}}D({\bm{x}}_{t})-{\bm{y}})$ through the decoder network and utilize the resulting gradient w.r.t.
-the latent as the error direction.
-We note that by setting ${\bm{e}}=-\partial C(\mathcal{D}({\bm{x}}_{t}))/\partial{\bm{x}}_{t}$ we only require backpropagation through the decoder model, which is significantly less expensive than backpropagating through the denoiser network, as done by previous works.
-For both degradations, we also include additive white Gaussian noise with $\sigma_{\bm{y}}=0.05$.
-We present the results on free-form inpainting and 8×8$\times$8 × super-resolution (SR) in Table 1.
-In inpainting, we find that our method generates better-aligned parts for the missing image regions, reflected in the significant improvement in FID, while obtaining similar PSNR and LPIPS scores to state-of-the-art methods.
-For super-resolution, although the improvements are smaller, we still find that we can compete with existing methods in both the quality and faithfulness of the generated images.
-The main advantage we showcase is inference time; our approach achieves similar or better results to every other existing method while only requiring a fraction of the computation.
-When we do not perform backpropagation through the VAE decoder (inpainting), our method requires only 15 seconds.
-In super-resolution, where we perform more optimization steps and backpropagate, the time increases to 1 minute.
-In contrast, P2L (Chung et al., 2024), the best-performing method, requires 30 minutes per single image since it relies on additional backpropagation steps through the denoiser to infer the text condition.
-Since there is no publicly available implementation for P2L we report the results directly from the paper and estimate the inference time as 5×5$\times$5 × that of LDPS (Chung et al., 2023), which the authors mentioned as a reasonable expectation as it linearly scales with the number of text prompt optimization iterations.
+For both degradations, we also include additive white Gaussian noise with $\sigma_{\bm{y}}=0.05$, use 20 DDIM [33] steps and normalize the computed gradient ${\bm{e}}_{t}$ with its $\infty$-norm.
+We present the results on free-form inpainting in Table 2.
+We find that our method generates better-aligned parts for the missing image regions, reflected in the significant improvement in FID.
+At the same time, our algorithm maintains consistency with the given image parts, which is reflected in the similar PSNR and LPIPS scores to the baselines.
+For $8\times$ super-resolution (Table 2), although the improvements are smaller, we attain similar quality and faithfulness to the generated images at a fraction of the inference time.
+The best-performing baseline, P2L [6], also utilizes a PaLI VLM [3] to caption the low-resolution images before the diffusion inference.
+We believe that the main advantage of P2L comes from introducing text conditioning, whereas all other methods rely only on the unconditional model.
+To verify this assumption, we run our algorithm for $\times 8$ super-resolution with text prompts generated from the downsampled images using Qwen-2.5 [2] as the VLM.
+The results clearly demonstrate that we can indeed bridge the performance gap to P2L, and even improve upon the PSNR and FID metrics when introducing text captions to the super-resolution process.
+Overall, the main advantages are in inference time and GPU memory.
+We achieve similar or better results to every other method while only requiring a fraction of the compute.
+When we do not backpropagate through the VAE decoder (inpainting), our method requires only 15 seconds.
+In super-resolution, where we run more optimization steps and backpropagate, the time increases to 1 minute.
+In comparison, P2L requires 30 minutes.
+Since P2L has no public implementation, we report the results directly from the paper and estimate the inference time as $5\times$ that of LDPS [5], which the authors mention as a reasonable expectation.
+Regarding memory, for a single image, our forward passes only require $\sim 9\text{GB}$ of memory, compared to backpropagation, which consumes $\sim 17\text{GB}$.
+Figure 3: Comparison between our method and existing algorithms on free-form inpainting and $8\times$ super-resolution.
+We directly use the images and results from [6] since there is no code available to replicate their method.
+Table 1: Quantitative evaluation (PSNR, LPIPS, FID) on free-form inpainting.
+Inpaint (Free-form)
+Method PSNR ↑ LPIPS ↓ FID ↓ Time
+P2L+ captions [6] 21.99 0.229 32.82 30m
+LDPS 21.54 0.332 46.72 6m
+PSLD [28] 20.92 0.251 40.57 8m
+Ours 21.73 0.258 19.39 15s
+Table 2: Quantitative evaluation (PSNR, LPIPS, FID) on $8\times$ super-resolution.
+Super-res ($\times 8$)
+Method PSNR ↑ LPIPS ↓ FID ↓ Time
+P2L+ captions [6] 23.38 0.386 51.81 30m
+LDPS 23.21 0.475 61.09 6m
+PSLD [28] 23.17 0.471 60.81 8m
+Ours 24.26 0.455 60.99 1m
+Ours+ captions 24.95 0.405 44.74 1m
 Box Inpainting
@@
 We further push the model by asking it to inpaint a box that covers 25% of the input image.
-In the comparisons we also include FreeDoM (Yu et al., 2023), which, although not previously shown on inpainting, claims to be a fast training-free inference approach for any condition.
-We also include the fine-tuned SD-Inpaint model (Rombach et al., 2022), which required 500k additional training steps.
+In the comparisons, we also include FreeDoM [44], which, although not previously shown on inpainting, claims to be a fast training-free inference approach for any condition.
+We also include the fine-tuned SD-Inpaint model [27], which required 500k additional training steps.
 We present the results on box inpainting in Table 4 and Figure 4.
@@
 Our method achieves both high quality and consistency in a shorter time as all previous methods backpropagate through the denoiser weights.
-In Appendix C.1 we ablate the steps $K$ and learning rate $\lambda$ hyperparameters.
+Figure 4: Qualitative evaluation of large area (box) inpainting on ImageNet.
+Our method achieves results closer to the fine-tuned inpainting model while requiring a fraction of the time to run per image compared to baselines.
+Number of steps and learning rate
+We ablate the number of optimization steps $K$ and learning rate $\lambda$ hyperparameters of our algorithm on the box inpainting task.
+The quantitative results are reported in Table 4, where we find that the original choice of $K=5$ steps and $\lambda=0.5$ achieve the best reconstruction and image quality metrics.
+In Appendix Figure 14, we present qualitative results where we observe that running fewer optimization steps gives blurrier results, which is expected as the known regions of the image also seem to not have converged to the given values.
+Using a higher learning rate leads to the model sometimes ’overshooting’ by inpainting the missing regions with realistic-looking parts that do not necessarily fit the rest of the image.
+Finite difference approximation
+In Appendix A.5 we ablate the finite difference step size $\delta$, and compare to the exact Jacobian-vector product computation.
+We find that our method is robust to the choice of $\delta$ and yields similar results to using the exact computation, while requiring less time.
+4.2 Non-linear Constraints
+Table 3: Quantitative evaluation on large area (box) inpainting.
+Our method outperforms all previous baselines and is the one closest to fine-tuning the diffusion model for inpainting (SD-Inpaint).
+When using fewer steps $K$, our algorithm does not sufficiently converge.
+When using a very high learning rate $\lambda$, it overshoots, adding non-realistic parts.
 Inpaint (Box)
 Method PSNR ↑ LPIPS ↓ FID ↓ Time
-LDPS 17.52 0.42 76.32 6min
-PSLD 17.30 0.38 74.02 8min
-FreeDoM 16.18 0.42 55.68 1min
-(Yu et al., 2023)
-Ours 18.30 0.30 42.01 15s
+LDPS 17.52 0.42 76.32 6m
+PSLD [28] 17.30 0.38 74.02 8m
+FreeDoM [44] 16.18 0.42 55.68 1m
+Ours ($K=5$, $\lambda=0.5$) 18.30 0.30 42.01 15s
+Ours ($K=2$, $\lambda=0.5$) 18.01 0.39 68.75 7s
+Ours ($K=5$, $\lambda=1.0$) 17.48 0.32 47.20 15s
 SD-Inpaint 19.05 0.28 32.93 4s
-Table 2: Quantitative evaluation (PSNR, LPIPS, FID) on large area (box) inpainting.
-4.2 Style-guided generation
-Our algorithm can be applied to non-linear differentiable constraints using backpropagation to approximate ${\bm{e}}$.
-Computing this gradient direction does not require backpropagation through the denoiser but through the separate, more lightweight ’critic’ network $f$.
-We demonstrate this with the non-linear constraint of style-guided generation.
+Table 4: Quantitative evaluation of style generation.
+The style score is what the gradient steps are directly optimizing for when using CLIP.
+Our approach achieves better style scores than the baselines without compromising faithfulness to the prompt, even when using a different model to guide style (OpenCLIP).
+Method Style Score ↓ CLIP ↑ Time
+DDIM 761.0 31.61 9s
+FreeDoM [44] 498.08 30.14 80s
+MPGD [13] 441.00 26.61 50s
+Ours ($w=2$) 368.37 23.95 45s
+Ours ($w=5$) 310.96 24.57 45s
+OursOpenCLIP ($w=5$) 434.45 25.94 45s
+Our algorithm can be applied to any non-linear differentiable constraint.
+The Newton derivation for non-linear constraints in Eq.
+(7) involves inverting two Jacobian matrices, the denoiser Jacobian $\bm{J}$} and the Jacobian of the constraint function $f$, ${\bm{J}}_{f}$.
+Our algorithm offers a way to approximate ${\bm{J}}^{-1}$.
+For ${\bm{J}}_{f}^{-1}$, we use the gradient descent direction ${\bm{J}}_{f}^{T}$, which is computed using backpropagation through the network $f$.
+Computing the gradient descent direction for $f$ does not require backpropagation through the denoiser, only through $f$.
+Thus, for non-linear constraints, we combine our proposed Newton direction for the denoiser with gradient descent for the differentiable constraint.
+In this section, we showcase style-guided generation with our algorithm.
+In Appendix B.1, we present results on mask-guided generation, where we show that our method outperforms the strongest baseline, MPGD [13].
+We discuss non-differentiable constraints in Appendix B.3.
+Style-guided generation
 The goal is to generate an image that simultaneously follows the style of a reference image ${\bm{x}}_{ref}$ and a given text prompt.
-Following previous works (Yu et al., 2023; He et al., 2024) which also perform style-guided generation, we do this by matching the statistics of CLIP (Radford et al., 2021) features between the reference and the generated images.
-We minimize the Frobenius norm between the Gram matrices (Gatys et al., 2016) of the 2nd layer CLIP features, making the cost to be minimized
-$\|\text{Gram}(\text{CLIP}_{2}({\bm{x}}_{ref}))-\text{Gram}(\text{CLIP}_{2}(\hat{{\bm{x}}}_{0}))\|_{F}^{2}.$ (15)
-Adhering to the evaluation of Yu et al (2023), we use 1000 random pairs of reference style images from WikiArt (Tan et al., 2019) and prompts from PartiPromtps (Yu et al., 2022).
+Following previous works [44, 13], which also perform style-guided generation, we match the statistics of CLIP [25] features between the reference and the generated images while denoising with a text prompt.
+We use the 2nd CLIP layer features to define the cost $C$ as the Frobenius norm of the the Gram matrix difference [11] $C=\lVert\text{Gram}(\text{CLIP}_{2}({\bm{x}}_{ref}))-\text{Gram}(\text{CLIP}_{2}(\hat{{\bm{x}}}_{0}))\rVert_{F}^{2}$.
+Figure 5: We guide the style of Stable Diffusion images with a CLIP (or OpenCLIP) model, using classifier-free guidance $w=5$.
+The images generated by our algorithm are closer to the reference style while maintaining faithfulness to the text prompt.
+Adhering to the evaluation of Yu et al [44], we use 1000 random pairs of reference style images from WikiArt [37] and prompts from PartiPromtps [43].
 We measure the CLIP similarity between the generated images and the text prompts to evaluate the faithfulness to the text condition, and the final difference between the Gram matrices to evaluate the faithfulness to the style reference (style score).
 We use the CLIP ViT-B/16 model for guiding the style of the image and evaluating.
-We also repeat the experiment using the OpenCLIP ViT-B/32 model (Cherti et al., 2023) for guidance.
-We perform $K=5$ gradient updates for every denoising step, using a linearly decreasing learning rate $\lambda$ from 0.5 to 0.1 and classifier-free guidance (Ho & Salimans, 2022) $w=2$ and $w=5$ for the denoiser.
+We also repeat the experiment using the OpenCLIP ViT-B/32 model [4] for guidance.
+We perform $K=5$ gradient updates for every denoising step, using a linearly decreasing learning rate $\lambda$ from 0.5 to 0.1 and classifier-free guidance [15] $w=2$ and $w=5$ for the denoiser.
 The results are presented in Table 4 and Figure 5.
-We find that our method achieves the overall best style score.
-Our approach is the best at minimizing the constraint which is also the the evaluation metric.
-Although in general, the lower the style score the more difficult it is to maintain high CLIP similarity with the prompt, we observe that our method still balances between the style and text, especially when we increase the guidance weight.
-We also used an OpenCLIP model to guide the style.
-With a CLIP model still as the evaluator, we achieve better style scores than the baselines that optimized directly with CLIP.
-Our proposed update reduces the target cost better without generating adversarial artifacts that trick the CLIP evaluation.
-Regarding inference time, our method is the fastest out of all baselines, excluding the non-style-guided DDIM.
-The baseline of MPGD (He et al., 2024), which is comparable in inference speed, does not backpropagate the style loss through the denoiser network, but only modifies the $\hat{{\bm{x}}}_{0}$ estimation at every denoising step.
-As we demonstrate in the supplementary, MPGD completely fails to propagate the constraint to distant pixels when the conditioning affects a specific region in the image.
-This makes it unusable in some constrained sampling settings, such as box inpainting.
-Figure 6: Style-guided generation.
-We show the results of guiding Stable Diffusion with a CLIP and an OpenCLIP model, using classifier-free guidance $w=5$.
-The images generated by our algorithm follow the reference style better than previous algorithms.
-Method Style Score* ↓ CLIP ↑ Time
-DDIM 761.0 31.61 9s
-FreeDoM 498.8 30.14 80s
-MPGD (He et al., 2024) 441.0 26.61 50s
-Ours (w=2) 368.37 23.95 45s
-Ours (w=5) 310.96 24.57 45s
-Ours (OpenCLIP, w=5) 434.45 25.94 45s
-Table 3: Quantitative evaluation of style generation. *The Style Score is what the gradient is optimizing.
-4.3 Mask-guided generation
-As an additional non-linear constraint, we guide the Stable Diffusion model with a separately trained face segmentation network.
-We employ an off-the-shelf model from huggingface and set the constraint $C$ to be the KL divergence between the per-pixel segmentation classes predicted for a reference image and the generated image.
-Using 100 images from the FFHQ (Karras et al., 2019) validation set, we run both our method and the MPGD (He et al., 2024) baseline.
-We chose MPGD because as the fastest baseline that works very well with ’dense’ constraints, i.e.
+Our method is best at minimizing the constraint, which is the style score evaluation metric.
+Although in general, the lower the style score the more difficult it is to maintain high CLIP similarity with the prompt, we observe that our method balances well between the style and text, especially when we increase the guidance weight.
+When using an OpenCLIP model to guide the style, we achieve a better style score than the baselines that optimized directly for it with CLIP.
+This also indicates that we minimize the target cost without generating adversarial artifacts that trick the CLIP evaluation.
+Excluding the non-guided DDIM, our method is as fast as MPGD [13], which does not backpropagate the style loss through the denoiser network, but only modifies the $\hat{{\bm{x}}}_{0}$ estimation at every denoising step.
+By only adjusting the $\hat{{\bm{x}}}_{0}$ prediction, MPGD completely fails to propagate the constraint to distant pixels (Appendix  B.7), making it unusable in other constrained sampling settings.
+Mask-guided generation
+We guide the Stable Diffusion model with a separately trained face segmentation network.
+We employ an off-the-shelf model and set the constraint $C$ to be the KL divergence between the per-pixel segmentation classes predicted for a reference image and the generated image.
+Using 100 images from the FFHQ [17] validation set, we run both our method and MPGD [13], which is the fastest baseline that works very well with ’dense’ constraints, i.e.
 constraints that are applied to all pixels.
+Table 5: Mask-conditioned generation using 100 FFHQ validation set images as reference, using Stable Diffusion with the prompt ’a headshot photo’.
+Method mIoU ↑ CLIP-FID ↓
+DDIM 0.09 48.78
+MPGD [13] ($\rho=1$) 0.47 77.11
+MPGD ($\rho=0.5$) 0.36 56.38
+Ours 0.42 59.79
 The results in Table 5 show that with a similar compute budget, our method achieves both faithfulness (mIoU between generated and reference images) and image quality (CLIP-FID).
-We used CLIP-FID because it performs better than Inception-FID on a small set of images (Kynkäänniemi et al., 2023).
-MPGD with a high weight ($\rho$) ’burns in’ the segmentation mask, leading to non-realistic images, whereas a lower weight does not produce images faithful to the mask.
-We provide qualitative examples of the generated images and the masks used in the Appendix C.5.
-Method mIoU ↑ CLIP-FID ↓ Time
-DDIM 0.09 48.78 3s
-MPGD ($\rho=1$) 0.47 77.11 1min
-MPGD ($\rho=0.5$) 0.36 56.38 1min
-Ours 0.42 59.79 1min
-Table 4: Mask-conditioned generation using 100 FFHQ validation set images as reference, using Stable Diffusion with the prompt ’a headshot photo’.
+We used CLIP-FID because it performs better than Inception-FID on a small set of images [18].
+MPGD with a high weight ($\rho$) ’burns in’ the segmentation mask, leading to artifacts and non-realistic images, whereas a lower weight does not produce images faithful to the mask.
+We provide qualitative results of our segmentation mask-guided generation experiment in Appendix Figure 13.
+The baseline (MPGD) either over-satisfies the constraint by burning in artifacts ($\rho=1$) or fails to generate images that adhere to the constraint ($\rho=0.5$).
+Our approach generates the most realistic images while also getting the mask prediction to match to the reference image.
+Using Stable Diffusion, we chose the text prompt ’a headshot photo’ to constrain the generated images.
+Considering the limitations of generating faces with Stable Diffusion, the results may not be on par with an FFHQ-specific model, but we still find our algorithm able to synthesize more usable images than the baseline, even in this difficult case.
+Limitations
+Using high classifier-free guidance ($w>10$), which is useful in some training-free tasks such as multi-view inference [42], requires us to significantly reduce the learning rate $\lambda$.
+The high guidance alters the spectral properties of the denoiser Jacobian, making the model more sensitive to small changes in the input.
+Another limitation we highlight is with distilled models [35, 21] that perform inference in fewer (1-5) steps.
+We observed that our algorithm requires more steps to achieve comparable results (Appendix B.4), mitigating the inference speed benefits of distilling the model.
 5 Conclusion
-In this work, we presented a new algorithm for fast inference in pre-trained diffusion models under constraints.
-Our approach exploits the relationship between the noisy and clean images during the denoising process to perform a gradient update that requires no expensive backpropagation operations through the diffusion model.
-Our method produces results comparable to the state-of-the-art with a fraction of the inference time.
-Our approach offers a practical method to sample from large pre-trained generative image models under any condition and has the potential to enable training-free downstream applications that rely on utilizing a strong image prior.
-Impact Statement
-This paper presents work whose goal is to advance the field of Machine Learning.
-We acknowledge that there are many potential societal consequences of our work.
-Beyond the positives, we highlight the ability to edit images with the intent to deceive and mislead as requiring careful consideration.
+We presented a new algorithm for inference under arbitrary constraints in pre-trained diffusion models.
+Our approach exploits the relationship between the noisy image input and clean image output of the denoiser to approximate the Newton optimization steps with cheap forward passes.
+The images generated under linear and non-linear constraints are comparable to or better than state-of-the-art methods, at a fraction of the inference time.
+We offer a practical algorithm to sample from large pre-trained generative image models under any condition, with the potential to enable new training-free downstream applications that rely on a strong image prior.
 Acknowledgments
-Part of this work was done during an internship at Microsoft Research, Redmond. This research was also partially supported by NSF grants IIS-2123920, IIS-2212046.
-Appendix A Analysis of the proposed algorithm
-A.1 Convergence
-We resort to inexact Newton methods (Dembo et al., 1982) to analyze the convergence and validity of the proposed approach. The original Newton’s method aims to solve the system
+Part of this work was done during an internship at Microsoft Research Redmond. This research was also partially supported by NSF grants IIS-2123920, IIS-2212046.
+NeurIPS Paper Checklist
+  1. 1.
+    Claims
+  2. Question: Do the main claims made in the abstract and introduction accurately reflect the paper’s contributions and scope?
+  3. Answer: [Yes]
+  4.
+Justification: In the abstract and introduction we clearly state our contribution of a new algorithm for constrained sampling in pre-trained diffusion models, and how it improves over existing baselines on a set of linear and non-linear tasks.
+We also refer to the mechanism of our algorithm, approximating the Newton steps, which is shown in the method section.
+  5. Guidelines:
+    • The answer NA means that the abstract and introduction do not include the claims made in the paper.
+    • The abstract and/or introduction should clearly state the claims made, including the contributions made in the paper and important assumptions and limitations.
+A No or NA answer to this question will not be perceived well by the reviewers.
+    • The claims made should match theoretical and experimental results, and reflect how much the results can be expected to generalize to other settings.
+    • It is fine to include aspirational goals as motivation as long as it is clear that these goals are not attained by the paper.
+  6. 2.
+    Limitations
+  7. Question: Does the paper discuss the limitations of the work performed by the authors?
+  8. Answer: [Yes]
+  9.
+Justification: We included a paragraph describing the limitations of our method concerning its applicability to tasks that require high classifier-free guidance and on models distilled from the base, large diffusion model.
+  10. Guidelines:
+    • The answer NA means that the paper has no limitation while the answer No means that the paper has limitations, but those are not discussed in the paper.
+    • The authors are encouraged to create a separate "Limitations" section in their paper.
+    • The paper should point out any strong assumptions and how robust the results are to violations of these assumptions (e.g., independence assumptions, noiseless settings, model well-specification, asymptotic approximations only holding locally).
+The authors should reflect on how these assumptions might be violated in practice and what the implications would be.
+    • The authors should reflect on the scope of the claims made, e.g., if the approach was only tested on a few datasets or with a few runs.
+In general, empirical results often depend on implicit assumptions, which should be articulated.
+    • The authors should reflect on the factors that influence the performance of the approach.
+For example, a facial recognition algorithm may perform poorly when image resolution is low or images are taken in low lighting.
+Or a speech-to-text system might not be used reliably to provide closed captions for online lectures because it fails to handle technical jargon.
+    • The authors should discuss the computational efficiency of the proposed algorithms and how they scale with dataset size.
+    • If applicable, the authors should discuss possible limitations of their approach to address problems of privacy and fairness.
+    • While the authors might fear that complete honesty about limitations might be used by reviewers as grounds for rejection, a worse outcome might be that reviewers discover limitations that aren’t acknowledged in the paper.
+The authors should use their best judgment and recognize that individual actions in favor of transparency play an important role in developing norms that preserve the integrity of the community.
+Reviewers will be specifically instructed to not penalize honesty concerning limitations.
+  11. 3.
+    Theory assumptions and proofs
+  12. Question: For each theoretical result, does the paper provide the full set of assumptions and a complete (and correct) proof?
+  13. Answer: [Yes]
+  14.
+Justification: Our main contribution is showing how we can approximate the Newton steps with a cheap and fast alternative.
+We show why this approximation is principled using the inexact Newton method in the main text and further discuss it in the appendix.
+  15. Guidelines:
+    • The answer NA means that the paper does not include theoretical results.
+    • All the theorems, formulas, and proofs in the paper should be numbered and cross-referenced.
+    • All assumptions should be clearly stated or referenced in the statement of any theorems.
+    • The proofs can either appear in the main paper or the supplemental material, but if they appear in the supplemental material, the authors are encouraged to provide a short proof sketch to provide intuition.
+    • Inversely, any informal proof provided in the core of the paper should be complemented by formal proofs provided in appendix or supplemental material.
+    • Theorems and Lemmas that the proof relies upon should be properly referenced.
+  16. 4.
+    Experimental result reproducibility
+  17.
+Question: Does the paper fully disclose all the information needed to reproduce the main experimental results of the paper to the extent that it affects the main claims and/or conclusions of the paper (regardless of whether the code and data are provided or not)?
+  18. Answer: [Yes]
+  19.
+Justification: Reproducing our algorithm is simple by following the steps we outline in Algorithm 1.
+We also include code in the supplementary material to showcase a simple implementation of our method using Stable Diffusion.
+  20. Guidelines:
+    • The answer NA means that the paper does not include experiments.
+    • If the paper includes experiments, a No answer to this question will not be perceived well by the reviewers: Making the paper reproducible is important, regardless of whether the code and data are provided or not.
+    • If the contribution is a dataset and/or model, the authors should describe the steps taken to make their results reproducible or verifiable.
+    • Depending on the contribution, reproducibility can be accomplished in various ways.
+For example, if the contribution is a novel architecture, describing the architecture fully might suffice, or if the contribution is a specific model and empirical evaluation, it may be necessary to either make it possible for others to replicate the model with the same dataset, or provide access to the model.
+In general.
+releasing code and data is often one good way to accomplish this, but reproducibility can also be provided via detailed instructions for how to replicate the results, access to a hosted model (e.g., in the case of a large language model), releasing of a model checkpoint, or other means that are appropriate to the research performed.
+    • While NeurIPS does not require releasing code, the conference does require all submissions to provide some reasonable avenue for reproducibility, which may depend on the nature of the contribution.
+For example
+      1. (a)
+        If the contribution is primarily a new algorithm, the paper should make it clear how to reproduce that algorithm.
+      2. (b)
+        If the contribution is primarily a new model architecture, the paper should describe the architecture clearly and fully.
+      3. (c)
+        If the contribution is a new model (e.g., a large language model), then there should either be a way to access this model for reproducing the results or a way to reproduce the model (e.g., with an open-source dataset or instructions for how to construct the dataset).
+      4. (d)
+        We recognize that reproducibility may be tricky in some cases, in which case authors are welcome to describe the particular way they provide for reproducibility.
+In the case of closed-source models, it may be that access to the model is limited in some way (e.g., to registered users), but it should be possible for other researchers to have some path to reproducing or verifying the results.
+  21. 5.
+    Open access to data and code
+  22. Question: Does the paper provide open access to the data and code, with sufficient instructions to faithfully reproduce the main experimental results, as described in supplemental material?
+  23. Answer: [Yes]
+  24.
+Justification: The models (Stable Diffusion) and datasets (ImageNet, WikiArt, PartiPrompts, FFHQ) used are all publicly available.
+Furthermore, we include the code to reproduce the algorithm in the supplementary material.
+  25. Guidelines:
+    • The answer NA means that paper does not include experiments requiring code.
+    • Please see the NeurIPS code and data submission guidelines (https://nips.cc/public/guides/CodeSubmissionPolicy) for more details.
+    • While we encourage the release of code and data, we understand that this might not be possible, so “No” is an acceptable answer.
+Papers cannot be rejected simply for not including code, unless this is central to the contribution (e.g., for a new open-source benchmark).
+    • The instructions should contain the exact command and environment needed to run to reproduce the results.
+See the NeurIPS code and data submission guidelines (https://nips.cc/public/guides/CodeSubmissionPolicy) for more details.
+    • The authors should provide instructions on data access and preparation, including how to access the raw data, preprocessed data, intermediate data, and generated data, etc.
+    • The authors should provide scripts to reproduce all experimental results for the new proposed method and baselines.
+If only a subset of experiments are reproducible, they should state which ones are omitted from the script and why.
+    • At submission time, to preserve anonymity, the authors should release anonymized versions (if applicable).
+    • Providing as much information as possible in supplemental material (appended to the paper) is recommended, but including URLs to data and code is permitted.
+  26. 6.
+    Experimental setting/details
+  27. Question: Does the paper specify all the training and test details (e.g., data splits, hyperparameters, how they were chosen, type of optimizer, etc.) necessary to understand the results?
+  28. Answer: [Yes]
+  29.
+Justification: We describe in detail each experiment we performed, including the exact data splits and how we sampled from them.
+We also describe in detail the hyperparameters chosen for our algorithm.
+  30. Guidelines:
+    • The answer NA means that the paper does not include experiments.
+    • The experimental setting should be presented in the core of the paper to a level of detail that is necessary to appreciate the results and make sense of them.
+    • The full details can be provided either with the code, in appendix, or as supplemental material.
+  31. 7.
+    Experiment statistical significance
+  32. Question: Does the paper report error bars suitably and correctly defined or other appropriate information about the statistical significance of the experiments?
+  33. Answer: [No]
+  34.
+Justification: It is not common in related works on training-free constrained sampling from generative models to report error bars on experiments regarding the generated image quality.
+Following the previous work, we only performed our experiments once.
+  35. Guidelines:
+    • The answer NA means that the paper does not include experiments.
+    • The authors should answer "Yes" if the results are accompanied by error bars, confidence intervals, or statistical significance tests, at least for the experiments that support the main claims of the paper.
+    • The factors of variability that the error bars are capturing should be clearly stated (for example, train/test split, initialization, random drawing of some parameter, or overall run with given experimental conditions).
+    • The method for calculating the error bars should be explained (closed form formula, call to a library function, bootstrap, etc.)
+    • The assumptions made should be given (e.g., Normally distributed errors).
+    • It should be clear whether the error bar is the standard deviation or the standard error of the mean.
+    • It is OK to report 1-sigma error bars, but one should state it.
+The authors should preferably report a 2-sigma error bar than state that they have a 96% CI, if the hypothesis of Normality of errors is not verified.
+    • For asymmetric distributions, the authors should be careful not to show in tables or figures symmetric error bars that would yield results that are out of range (e.g. negative error rates).
+    • If error bars are reported in tables or plots, The authors should explain in the text how they were calculated and reference the corresponding figures or tables in the text.
+  36. 8.
+    Experiments compute resources
+  37. Question: For each experiment, does the paper provide sufficient information on the computer resources (type of compute workers, memory, time of execution) needed to reproduce the experiments?
+  38. Answer: [Yes]
+  39. Justification: We provide the compute setup we used and describe the inference time and memory requirements for running our and previous algorithms in the experiments section.
+  40. Guidelines:
+    • The answer NA means that the paper does not include experiments.
+    • The paper should indicate the type of compute workers CPU or GPU, internal cluster, or cloud provider, including relevant memory and storage.
+    • The paper should provide the amount of compute required for each of the individual experimental runs as well as estimate the total compute.
+    • The paper should disclose whether the full research project required more compute than the experiments reported in the paper (e.g., preliminary or failed experiments that didn’t make it into the paper).
+  41. 9.
+    Code of ethics
+  42. Question: Does the research conducted in the paper conform, in every respect, with the NeurIPS Code of Ethics https://neurips.cc/public/EthicsGuidelines?
+  43. Answer: [Yes]
+  44. Justification: There were no human subjects in our tests and all datasets used are publicly available.
+  45. Guidelines:
+    • The answer NA means that the authors have not reviewed the NeurIPS Code of Ethics.
+    • If the authors answer No, they should explain the special circumstances that require a deviation from the Code of Ethics.
+    • The authors should make sure to preserve anonymity (e.g., if there is a special consideration due to laws or regulations in their jurisdiction).
+  46. 10.
+    Broader impacts
+  47. Question: Does the paper discuss both potential positive societal impacts and negative societal impacts of the work performed?
+  48. Answer: [Yes]
+  49. Justification: We included a paragraph discussing the societal impacts of our work in the appendix.
+  50. Guidelines:
+    • The answer NA means that there is no societal impact of the work performed.
+    • If the authors answer NA or No, they should explain why their work has no societal impact or why the paper does not address societal impact.
+    • Examples of negative societal impacts include potential malicious or unintended uses (e.g., disinformation, generating fake profiles, surveillance), fairness considerations (e.g., deployment of technologies that could make decisions that unfairly impact specific groups), privacy considerations, and security considerations.
+    • The conference expects that many papers will be foundational research and not tied to particular applications, let alone deployments.
+However, if there is a direct path to any negative applications, the authors should point it out.
+For example, it is legitimate to point out that an improvement in the quality of generative models could be used to generate deepfakes for disinformation.
+On the other hand, it is not needed to point out that a generic algorithm for optimizing neural networks could enable people to train models that generate Deepfakes faster.
+    • The authors should consider possible harms that could arise when the technology is being used as intended and functioning correctly, harms that could arise when the technology is being used as intended but gives incorrect results, and harms following from (intentional or unintentional) misuse of the technology.
+    • If there are negative societal impacts, the authors could also discuss possible mitigation strategies (e.g., gated release of models, providing defenses in addition to attacks, mechanisms for monitoring misuse, mechanisms to monitor how a system learns from feedback over time, improving the efficiency and accessibility of ML).
+  51. 11.
+    Safeguards
+  52.
+Question: Does the paper describe safeguards that have been put in place for responsible release of data or models that have a high risk for misuse (e.g., pretrained language models, image generators, or scraped datasets)?
+  53. Answer: [N/A]
+  54. Justification: We release no new models or datasets. The existing models we utilize (Stable Diffusion) include such safeguards.
+  55. Guidelines:
+    • The answer NA means that the paper poses no such risks.
+    • Released models that have a high risk for misuse or dual-use should be released with necessary safeguards to allow for controlled use of the model, for example by requiring that users adhere to usage guidelines or restrictions to access the model or implementing safety filters.
+    • Datasets that have been scraped from the Internet could pose safety risks. The authors should describe how they avoided releasing unsafe images.
+    • We recognize that providing effective safeguards is challenging, and many papers do not require this, but we encourage authors to take this into account and make a best faith effort.
+  56. 12.
+    Licenses for existing assets
+  57.
+Question: Are the creators or original owners of assets (e.g., code, data, models), used in the paper, properly credited and are the license and terms of use explicitly mentioned and properly respected?
+  58. Answer: [Yes]
+  59. Justification: All models and datasets we used are publicly available and open for research use.
+  60. Guidelines:
+    • The answer NA means that the paper does not use existing assets.
+    • The authors should cite the original paper that produced the code package or dataset.
+    • The authors should state which version of the asset is used and, if possible, include a URL.
+    • The name of the license (e.g., CC-BY 4.0) should be included for each asset.
+    • For scraped data from a particular source (e.g., website), the copyright and terms of service of that source should be provided.
+    • If assets are released, the license, copyright information, and terms of use in the package should be provided.
+For popular datasets, paperswithcode.com/datasets has curated licenses for some datasets.
+Their licensing guide can help determine the license of a dataset.
+    • For existing datasets that are re-packaged, both the original license and the license of the derived asset (if it has changed) should be provided.
+    • If this information is not available online, the authors are encouraged to reach out to the asset’s creators.
+  61. 13.
+    New assets
+  62. Question: Are new assets introduced in the paper well documented and is the documentation provided alongside the assets?
+  63. Answer: [N/A]
+  64. Justification: There are no new assets introduced in the paper.
+  65. Guidelines:
+    • The answer NA means that the paper does not release new assets.
+    • Researchers should communicate the details of the dataset/code/model as part of their submissions via structured templates. This includes details about training, license, limitations, etc.
+    • The paper should discuss whether and how consent was obtained from people whose asset is used.
+    • At submission time, remember to anonymize your assets (if applicable). You can either create an anonymized URL or include an anonymized zip file.
+  66. 14.
+    Crowdsourcing and research with human subjects
+  67.
+Question: For crowdsourcing experiments and research with human subjects, does the paper include the full text of instructions given to participants and screenshots, if applicable, as well as details about compensation (if any)?
+  68. Answer: [N/A]
+  69. Justification: No human subject experiments were performed.
+  70. Guidelines:
+    • The answer NA means that the paper does not involve crowdsourcing nor research with human subjects.
+    • Including this information in the supplemental material is fine, but if the main contribution of the paper involves human subjects, then as much detail as possible should be included in the main paper.
+    • According to the NeurIPS Code of Ethics, workers involved in data collection, curation, or other labor should be paid at least the minimum wage in the country of the data collector.
+  71. 15.
+    Institutional review board (IRB) approvals or equivalent for research with human subjects
+  72.
+Question: Does the paper describe potential risks incurred by study participants, whether such risks were disclosed to the subjects, and whether Institutional Review Board (IRB) approvals (or an equivalent approval/review based on the requirements of your country or institution) were obtained?
+  73. Answer: [N/A]
+  74. Justification: No human subject experiments were performed.
+  75. Guidelines:
+    • The answer NA means that the paper does not involve crowdsourcing nor research with human subjects.
+    • Depending on the country in which research is conducted, IRB approval (or equivalent) may be required for any human subjects research.
+If you obtained IRB approval, you should clearly state this in the paper.
+    • We recognize that the procedures for this may vary significantly between institutions and locations, and we expect authors to adhere to the NeurIPS Code of Ethics and the guidelines for their institution.
+    • For initial submissions, do not include any information that would break anonymity (if applicable), such as the institution conducting the review.
+  76. 16.
+    Declaration of LLM usage
+  77.
+Question: Does the paper describe the usage of LLMs if it is an important, original, or non-standard component of the core methods in this research?
+Note that if the LLM is used only for writing, editing, or formatting purposes and does not impact the core methodology, scientific rigorousness, or originality of the research, declaration is not required.
+  78. Answer: [N/A]
+  79. Justification: No LLMs were used.
+  80. Guidelines:
+    • The answer NA means that the core method development in this research does not involve LLMs as any important, original, or non-standard components.
+    • Please refer to our LLM policy (https://neurips.cc/Conferences/2025/LLM) for what should or should not be described.
+Appendix A Further analysis of the proposed algorithm
+A.1 Spectra of the denoiser Jacobian
+In the main text, we resorted to inexact Newton methods [7] to analyze the convergence of the proposed algorithm.
+We reiterate that the original Newton (or Gauss-Newton) method aims to solve the system
 ${\bm{J}}{\bm{e}}_{t}={\bm{e}}.$ (16)
-In cases where computing the inverse of the Jacobian, as is the case of the diffusion denoiser Jacobian, inexact Newton methods propose first finding an approximation ${\bm{e}}_{t}^{*}$ to the solution of Eq (16) and then performing the Newton step using the approximate solution and reiterating until convergence.
-For inexact Newton methods to converge to the correct solution, the residual must be strictly reduced at every step
+For a general discussion on how to solve this system of equations we refer the reader to Gavin [12].
+In our case of denoising diffusion, where computing the inverse of $\bm{J}$} is expensive, inexact Newton methods utilize an approximate solution ${\bm{e}}_{t}^{*}$ to Eq (16), which is guaranteed to converge if the residual is strictly reduced at every step
 ${\bm{r}}={\bm{e}}-{\bm{J}}{\bm{e}}_{t}^{*},\quad\frac{\lVert{\bm{r}}\rVert_{2}}{\lVert{\bm{e}}\rVert_{2}}<\eta,\ \eta\in[0,1).$ (17)
-If we substitute the proposed update ${\bm{e}}_{t}^{*}=\lambda{\bm{J}}{\bm{e}}$ (Eq. 10) we get the residual
-${\bm{r}}={\bm{e}}-\lambda{\bm{J}}^{2}{\bm{e}}=({\bm{I}}-\lambda{\bm{J}}^{2}){\bm{e}}.$ (18)
-For this residual to be strictly reduced at every iteration of our algorithm, we need
+When substituting the proposed update ${\bm{e}}_{t}^{*}=\lambda{\bm{J}}{\bm{e}}$ (Eq. 10) we get
 $\frac{\lVert{\bm{r}}\rVert_{2}}{\lVert{\bm{e}}\rVert_{2}}=\frac{\lVert({\bm{I}}-\lambda{\bm{J}}^{2}){\bm{e}}\rVert_{2}}{\lVert{\bm{e}}\rVert_{2}}\leq\frac{\lVert{\bm{I}}-\lambda{\bm{J}}^{2}\rVert_{2}\lVert{\bm{e}}\rVert_{2}}{\lVert{\bm{e}}\rVert_{2}}=\lVert{\bm{I}}-\lambda{\bm{J}}^{2}\rVert_{2}<\eta.$ (18)
-Therefore, we need to show that the spectral norm of the matrix ${\bm{I}}-\lambda{\bm{J}}^{2}$ is smaller than 1 for a correct choice of learning rate $\lambda$.
-If ${\bm{J}}$ was diagonalizable (or more generally normal) then we could directly estimate $\lVert{\bm{I}}-\lambda{\bm{J}}^{2}\rVert_{2}$ using the largest eigenvalue of ${\bm{J}}$, following the spectral theorem.
-While we cannot directly assume that ${\bm{J}}$ is normal, we know that ${\bm{J}}$ should be ’almost’ symmetric, as is the optimal denoiser solution shown in Eq.
-14.
-This means that we could express ${\bm{J}}={\bm{S}}+{\bm{K}}$, where ${\bm{S}}=({\bm{J}}+{\bm{J}}^{T})/2$ is the symmetric component and ${\bm{K}}=({\bm{J}}-{\bm{J}}^{T})/2$ the skew-symmetric component, and expect $\lVert{\bm{K}}\rVert\ll\lVert{\bm{S}}\rVert$.
-In that case, we can approximate $\lVert{\bm{I}}-\lambda{\bm{J}}^{2}\rVert_{2}$ as
+Therefore, we need to show that the spectral norm of the matrix ${\bm{I}}-\lambda{\bm{J}}^{2}$ is strictly less than 1 for a correct choice of learning rate $\lambda$.
+If $\bm{J}$} was diagonalizable (or generally normal), then we could directly estimate $\lVert{\bm{I}}-\lambda{\bm{J}}^{2}\rVert_{2}$ using the largest eigenvalue of $\bm{J}$}.
+While we cannot directly assume that $\bm{J}$} is normal (we explicitly mentioned that $\bm{J}$} is not guaranteed to be symmetric), we know that $\bm{J}$} should be ’almost’ symmetric, as is the optimal denoiser solution shown in Eq.
+(15).
+This means that we could express ${\bm{J}}={\bm{S}}+{\bm{K}}$, where ${\bm{S}}=({\bm{J}}+{\bm{J}}^{T})/2$ is the symmetric and ${\bm{K}}=({\bm{J}}-{\bm{J}}^{T})/2$ the skew-symmetric component.
+Since both the symmetric and skew-symmetric components are normal we can estimate their spectral norms.
+In the case where $\lVert{\bm{K}}\rVert_{2}\ll\lVert{\bm{S}}\rVert_{2}$, which we expect since $\bm{J}$} is ’almost’ symmetric, we can approximate $\lVert{\bm{I}}-\lambda{\bm{J}}^{2}\rVert_{2}$ as
 $\lVert{\bm{I}}-\lambda{\bm{J}}^{2}\rVert_{2}\approx\lvert 1-\lambda\max_{i}\mu_{i}^{2}\rvert$ (19)
 where $\mu_{i}$ are the eigenvalues of $\bm{S}$}.
-To get an estimate of the magnitudes of the largest eigenvalues of ${\bm{S}}$, which are real, and of the eigenvalues of ${\bm{K}}$, which are imaginary, we use the Arnoldi iteration (Arnoldi, 1951).
+To get an estimate of the magnitudes of the largest eigenvalues of $\bm{S}$}, which are real, and of the eigenvalues of $\bm{K}$}, which are imaginary, we use the Arnoldi iteration [1].
 For the Arnoldi iteration we only require access to the matrix-vector products ${\bm{J}}{\bm{v}}$ and ${\bm{J}}^{T}{\bm{v}}$, which can be computed using the approximation of Eq.
@@
 We see that the contribution of the symmetric part of the matrix is the strongest, validating our assumption that the diffusion model’s Jacobian $\bm{J}$} is ’almost’ symmetric.
-If we approximate the spectral norm of Eq.( 19) using the symmetric component of the Jacobian we see that for a correct choice of $\lambda$ we can ensure that $\lvert 1-\lambda\max_{i}\mu_{i}\rvert<1$.
+If we approximate the spectral norm of Eq.( 18) using the symmetric component of the Jacobian we see that for a correct choice of $\lambda$ we can ensure that $\lvert 1-\lambda\max_{i}\mu_{i}\rvert<1$.
 Now we revisit the example of Fig.
@@
 These initial oscillations may be important for the high-quality results, as using lower learning rate did not attain similar-quality images.
-A.2 Difference between ${\bm{J}}$ and ${\bm{J}}^{T}$
+A.2 Symmetry of the Jacobian and difference in updates
+Symmetry
+Although, theoretically, the Jacobian of the denoiser should be symmetric (Eq.
+15), the trained diffusion model does not exactly match the real score and can yield non-symmetric Jacobians.
+In Figure 7, we perform a simple experiment to visualize this difference; we select a random image from the ImageNet [8] validation set and employ the Stable Diffusion 1.5 model [23] to denoise at three different noise levels $t=\{900,700,400\}$.
+We encode the image using the VAE encoder, scale it and add appropriate noise to get the intermediate diffusion latent for each timestep.
+We then give the noisy image to the denoiser network and compute the gradients $\partial\hat{{\bm{x}}}_{0}^{k,l}/\partial{\bm{x}}_{t}^{i,j}$ and $\partial\hat{{\bm{x}}}_{0}^{i,j}/\partial{\bm{x}}_{t}^{k,l}$ for randomly chosen pixels $(i,j),(k,l)$ using backpropagation.
+When we plot the gradients and compute the correlation coefficient $r$ we observe that the values deviate from $y=x$, which would indicate a symmetric Jacobian.
+Figure 7: Sample pairs $(i,j),(k,l)$ of the denoiser Jacobian $\nabla_{{\bm{x}}_{t}}\hat{{\bm{x}}_{0}}=\nabla_{{\bm{x}}_{t}}E[{\bm{x}}_{0}|{\bm{x}}_{t}]$ for different timesteps $t$.
+We observe that the Jacobian is not symmetric, justifying the difference between the proposed update steps and the previously used gradient descent steps.
+Figure 8: We present a simple experiment to compare the proposed gradient update ${\bm{J}}{\bm{e}}$ with that of previous methods ${\bm{J}}^{T}{\bm{e}}$.
+The goal is to update a noisy version of the image shown in (a) at $t=800$, such that the diffusion model’s prediction of the final image includes a red square in the middle, shown in (b).
+The results of gradient descent in (c) and our gradient update in (d) demonstrate that updates performed by previous methods lead to a considerably different result than the update we propose in this paper.
+Toy experiment
+Having shown that the Jacobian is not symmetric, we expect to find differences between the gradient updates of gradient descent (Eq.
+2.2) and our proposed update (Eq.
+10).
+To highlight these differences, we set up a toy experiment where we update an identical initial image with the two different directions, under the same condition.
+The experiment is showcased in Figure 8 and again utilizes the Stable Diffusion 1.5 model.
+We create a synthetic black-and-white grid image (Figure 8 (a)), which we blur, deterministically encode, scale and noise to get a diffusion latent at $t=800$.
+We then set up a constraint where we add a red square to the center of the original image (Figure 8 (b)).
+Instead of decoding the image and applying the constraint in image space, we also encode the constraint image and apply it directly in the Stable Diffusion latent space.
+We run 5 gradient updates with the same learning rate of $\lambda=1$ for each and demonstrate the final predicted $\hat{{\bm{x}}}_{0}$ for the ${\bm{J}}^{T}{\bm{e}}$ update of Eq.
+(2.2) (a) (Figure 8 (c)) and the proposed ${\bm{J}}{\bm{e}}$ of Eq (10) (Figure 8 (d)).
+The final result shows how the model intends to change the entire image when asked to add a red square in the middle.
+The resulting images differ substantially, with the proposed direction producing a more coherent image that tries to copy the newly introduced texture to the correct locations, i.e.
+the intersections of the lines.
+Although this is an empirical observation, we hypothesize that the two different directions can have vastly different effects on the image.
+In Figure 9 we repeat this experiment over multiple random seeds.
+Figure 9: Further runs of the toy experiment of A.2.
+More visualizations
 In the main text, we referred to the difference between $\bm{J}$} and ${\bm{J}}^{T}$ as a motivating factor for our approach and measured that difference by comparing elements $(i,j)$ and $(j,i)$ of the Jacobian matrix (Figure 7).
-In Figure 8 we further visualize the difference between using the Jacobian and its transpose by numerically computing how the Jacobian and the Jacobian transpose of the entire image would look for a single pixel, i.e.
+In Figure 10 we extend Fig.
+2.
+We visualize the difference between using the Jacobian and its transpose by numerically computing how the Jacobian and the Jacobian transpose of the entire image would look for a single pixel, i.e.
 which parts of the image are affected by a change in a single pixel.
@@
 ${\bm{J}}={\bm{J}}^{T}$, then a change in a single pixel would have a similar effect on the rest of the image, up to some noise because of the finite difference approximation.
-Contrary to the expectations, we find a significant difference between our proposed direction and backpropagation, with our approach having a better effect on retaining shapes and symmetry across the image.
-We see that in many cases, the model is trying to change ’corresponding’ parts of the image together, i.e.
-change both eyes or ears simultaneously, giving us a better understanding of the knowledge that the model has acquired regarding the image space in general.
-Appendix B Proposed algorithm as modification of DDIM
-To apply the proposed method we modified the DDIM sampling algorithm (Song et al., 2020).
-We provide a side-by-side comparison to show the difference between the original DDIM and the proposed algorithm.
-Extending our algorithm to other sampling methods should be intuitive by alternating between gradient updates from our algorithm and the diffusion updates computed with the diffusion sampling algorithm used.
-Algorithm 2 DDIM sampling (Song et al., 2020) using a pre-trained diffusion model.
-1:  Input: Pre-trained diffusion model $\hat{{\bm{x}}}_{0}({\bm{x}}_{t})$, diffusion step size $s$ and schedule parameters $\alpha_{i},\beta_{i}$
-2:  ${\bm{x}}_{T}\sim N(\textbf{0},{\bm{I}})$
-3:  for $t\in\{T,T-s,T-2s,\dots,s\}$ do
-4:     ${\bm{z}}_{t}\sim N(\textbf{0},{\bm{I}})$
-5:     $\mathbf{\epsilon}_{t}=\frac{1}{\sqrt{1-\alpha_{t-s}}}{\bm{x}}_{t}-\frac{\sqrt{\alpha_{t-s}}}{\sqrt{1-\alpha_{t-s}}}\hat{{\bm{x}}}_{0}({\bm{x}}_{t})$
-6:     ${\bm{x}}_{t-s}=\sqrt{\alpha_{t-s}}\hat{{\bm{x}}}_{0}({\bm{x}}_{t})+\sqrt{1-\alpha_{t-s}}\mathbf{\epsilon}_{t}+\beta_{t-s}{\bm{z}}_{t}$
-7:  end for
-8:  Return: ${\bm{x}}_{0}$
-Algorithm 3 The proposed algorithm for constrained sampling using a pre-trained diffusion model.
-1:  Input: Pre-trained diffusion model $\hat{{\bm{x}}}_{0}({\bm{x}}_{t})$, constraint $C({\bm{x}}_{0},{\bm{y}})=\lVert f(\hat{{\bm{x}}}_{0}({\bm{x}}_{t})-y\rVert_{2}^{2}$, condition ${\bm{y}}$, gradient step size $\delta$, iterations $K$, learning rate $\lambda$, diffusion step size $s$ and schedule parameters $\alpha_{i},\beta_{i}$
-2:  ${\bm{x}}_{T}\sim N(\textbf{0},{\bm{I}})$
-3:  for $t\in\{T,T-s,T-2s,\dots,s\}$ do
-4:     for $i\in\{1,2,\dots,K\}$ do
-5:        ${\bm{e}}={\bm{J}}_{f}^{-1}(f(\hat{{\bm{x}}}_{0}({\bm{x}}_{t}))-{\bm{y}})$Task-dependent}
-6:        ${\bm{e}}_{t}=[\hat{{\bm{x}}}_{0}({\bm{x}}_{t}+\delta{\bm{e}})-\hat{{\bm{x}}}_{0}({\bm{x}}_{t})]/\delta$
-7:        ${\bm{x}}_{t}={\bm{x}}_{t}+\lambda{\bm{e}}_{t}$
-8:     end for
-9:     ${\bm{z}}_{t}\sim N(\textbf{0},{\bm{I}})$
-10:     $\mathbf{\epsilon}_{t}=\frac{1}{\sqrt{1-\alpha_{t-s}}}{\bm{x}}_{t}-\frac{\sqrt{\alpha_{t-s}}}{\sqrt{1-\alpha_{t-s}}}\hat{{\bm{x}}}_{0}({\bm{x}}_{t})$
-11:     ${\bm{x}}_{t-s}=\sqrt{\alpha_{t-s}}\hat{{\bm{x}}}_{0}({\bm{x}}_{t})+\sqrt{1-\alpha_{t-s}}\mathbf{\epsilon}_{t}+\beta_{t-s}{\bm{z}}_{t}$
-12:  end for
-13:  Return: ${\bm{x}}_{0}$
+Contrary to expectations, we find a significant difference between our proposed direction and backpropagation, with our approach having a better effect on retaining shapes and symmetry across the image.
+We see that in many cases, the model is trying to change correlated parts of the image together, i.e.
+change both eyes or ears simultaneously, showing us some of the knowledge that the model has acquired regarding the image space in general through its text-to-image training.
 Figure 10: We visualize the difference between $\bm{J}$} and ${\bm{J}}^{T}$ by computing how the matrix wants to change the entire image would look for a perturbation of a single pixel.
@@
 Even though we use a numerical approximation, the proposed direction is sharper in regions, like the outline of the cat.
-Appendix C Additional Results
-C.1 Number of steps and learning rate ablation
+A.3 Inexact and exact Newton
+We set up a simple experiment on MNIST to compare the proposed inexact Newton update step with the exact Newton, i.e.
+computing the inverse of the Jacobian.
+We train a 25M parameter diffusion model, following the architecture of [9], on $32\times 32$ zero-padded MNIST images.
+For this model we can use auto-differentiation to compute the $1024\times 1024$ Jacobian matrix, which takes  10 seconds on our GPUs.
+We randomly sample images from the training dataset, add noise corresponding to $t=700$ and $t=500$ and denoise, predicting the $\hat{{\bm{x}}}_{0}$.
+Then, we aim to apply a simple edit to the predicted $\hat{{\bm{x}}}_{0}$ that increases or decreases the value of a single pixel in the image.
+For that edit, the error $\bm{e}$} corresponds to a +1 or -1 in the location of the edit.
+For each specific edit, we compute the update we should make to the noisy ${\bm{x}}_{t}$ using our inexact Newton method ${\bm{J}}{\bm{e}}$, the exact Newton ${\bm{J}}^{-1}{\bm{e}}$ and gradient descent ${\bm{J}}^{T}{\bm{e}}$.
+For the exact Newton method the Jacobian can be ill-conditioned, requiring us to utilize a pseudo-inverse in computing the update ${\bm{J}}^{-1}{\bm{e}}$.
+In Figure 11 we showcase the results of editing the image with the three update directions.
+The exact Newton step makes definite steps in updating the image, which converge faster to the desired solution.
+Our inexact step, qualitatively attempts to make similar edits to the image; e.g.
+when increasing the pixel intensity inside the digit ’0’, both updates also delete parts of the side, turning the digit into a 5 (Figure 11, left).
+Of course, as discussed the main paper, our inexact step requires multiple, smaller steps to achieve the result, but is much cheaper to compute.
+Finally, the gradient descent direction gives similar results to our proposed update, but as discussed in A.2, showcases qualitative differences that we hypothesize arise from the imperfect training of the diffusion model.
+Figure 11: We show the difference between the inexact Newton, exact Newton and gradient descent updates on a simple setting of generating MNIST digits.
+We simulate a constraint by choosing to increase or decrease the intensity of a single pixel in the image.
+Qualitatively, the updates made with the inexact and exact Newton methods are similar, with the inexact approach requiring much less compute and being easy to compute in practical settings.
+A.4 Inexact Newton steps in VAE space
+Figure 12: Comaprison between using backpropagation and Newton steps with linear constraints in image space.
+In Section 4.2, we described how instead of following the Newton recipe, which requires the inverse of the constraint Jacobian ${\bm{J}}_{f}^{-1}$, we used the gradient descent direction ${\bm{J}}_{f}^{T}$ when dealing with non-linear constraints.
+Here, we discuss a special case of non-linear constraints, constraints that are linear in image space.
+These constraints are still non-linear for latent diffusion models [27] since they involve the VAE decoder, which non-linearly transforms the latent representations that the diffusion model generates to RGB images.
+However, when the constraint is linear in image space we find that we can still utilize an inexact Newton approach to avoid backpropagation through the decoder. More specifically, for the case where
+$C({\bm{x}}_{t})=({\bm{A}}\mathcal{D}(\hat{{\bm{x}}}_{0}({\bm{x}}_{t}))-{\bm{y}})^{T}({\bm{A}}\mathcal{D}(\hat{{\bm{x}}}_{0}({\bm{x}}_{t}))-{\bm{y}})$ (20)
+we write the first-order Taylor approximation
+$C({\bm{x}}_{t}-{\bm{e}}_{t})\approx{\bm{A}}\mathcal{D}(\hat{{\bm{x}}}_{0}({\bm{x}}_{t}))-{\bm{A}}{\bm{J}}_{\mathcal{D}}{\bm{J}}{\bm{e}}_{t}-{\bm{y}})^{T}({\bm{A}}\mathcal{D}(\hat{{\bm{x}}}_{0}({\bm{x}}_{t}))-{\bm{A}}{\bm{J}}_{\mathcal{D}}{\bm{J}}{\bm{e}}_{t}-{\bm{y}})^{T}.$ (21)
+When we solve for $\nabla_{{\bm{e}}_{t}}C=0$ we get
+${\bm{J}}^{T}{\bm{J}}_{\mathcal{D}}^{T}{\bm{A}}^{T}({\bm{A}}\mathcal{D}(\hat{{\bm{x}}}_{0}({\bm{x}}_{t}))-{\bm{y}})={\bm{J}}^{T}{\bm{J}}_{\mathcal{D}}^{T}{\bm{J}}_{\mathcal{D}}{\bm{J}}{\bm{e}}_{t}.$ (22)
+Similarly to the results in the main paper, assuming that inverses exist, we end up with the following system
+${\bm{A}}^{T}({\bm{A}}\mathcal{D}(\hat{{\bm{x}}}_{0}({\bm{x}}_{t}))-{\bm{y}})={\bm{J}}_{\mathcal{D}}{\bm{J}}{\bm{e}}_{t}$ (23)
+which we rewrite as
+${\bm{e}}_{i}={\bm{J}}_{\mathcal{D}}{\bm{J}}{\bm{e}}_{t},\quad{\bm{e}}_{i}={\bm{A}}^{T}({\bm{A}}\mathcal{D}(\hat{{\bm{x}}}_{0}({\bm{x}}_{t}))-{\bm{y}}).$ (24)
+To apply the proposed Newton approach and get an update direction for ${\bm{x}}_{t}$, we must first solve the system ${\bm{e}}_{i}={\bm{J}}_{\mathcal{D}}{\bm{b}}$ for $\bm{b}$}, and then use the approximate solution ${\bm{e}}_{t}={\bm{J}}{\bm{b}}$.
+In the super-resolution experiments we performed in the paper, we opted for the gradient descent approach, which can be expressed as ${\bm{b}}={\bm{J}}_{\mathcal{D}}^{T}{\bm{e}}_{i}$.
+However, this requires backpropagating through the decoder model which in some cases may be inefficient or altogether unavailable.
+As an alternative, we can again resort to inexact Newton and use an ’approximate’ inverse to ${\bm{J}}_{\mathcal{D}}$, the encoder Jacobian $\bm{J}$$_{\mathcal{E}}$.
+Intuitively, the encoder model performs the inverse operation of the decoder, and therefore we could employ it to ’invert’ the decoding operation.
+Using the encoder allows us to replace backpropagation with forward passes by
+${\bm{b}}={\bm{J}}_{\mathcal{E}}{\bm{e}}_{i}\approx\left[\mathcal{E}(\mathcal{D}(\hat{{\bm{x}}}_{0})+\delta{\bm{e}}_{i})-\mathcal{E}(\mathcal{D}(\hat{{\bm{x}}}_{0}))\right]/\delta.$ (25)
+By combining the Newton step for the VAE space and the Newton step for the denoiser we can run our inference algorithm with no backpropagation operations.
+In Figure 12 we provide some qualitative comparisons of super-resolution with a Stable Diffusion model when using backpropagation through the decoder and the inexact Newton step in VAE space.
+Nevertheless, we opted for backpropagation in our experiments since the time required for multiple forward passes through the encoder and decoder models ended up being the same as backpropagating once through the decoder, and the memory requirements were not prohibitive.
+In cases where memory is an issue, the ’pure’ Newton approach could be an appealing alternative to avoid backpropagating through the decoder model.
+Using the the Jacobian of the encoder as an approximation for the inverse Jacobian of the decoder, has been discussed before in the context of autoencoders in Sorrenson et al.
+[36].
+To our knowledge, we are the first to employ this approximation for the diffusion autoencoders.
+A.5 Finite-difference approximation and exact gradient computation
+To compute the proposed update ${\bm{J}}{\bm{e}}$, we use the finite difference approximation ${\bm{J}}{\bm{e}}\approx(f({\bm{x}}+\delta{\bm{e}})-f({\bm{x}}))/\delta$.
+In comparison, the gradient descent direction ${\bm{J}}^{T}e$ is exactly computed using automatic differentiation, usually implemented as the backward gradient computation, e.g.
+e.backward() in PyTorch.
+Some libraries also offer forward-mode auto-differentiation, which directly computes the Jacobian-vector product ${\bm{J}}{\bm{e}}$.
+However, forward differentiation is not always implemented or optimized as well as the backward propagation of gradients.
+In the case of PyTorch, which is what we use for running our experiments, forward mode differentiation is not directly implemented for many of the custom layers of the Stable Diffusion model.
+To perform a comparison between our finite difference approximation and the exact forward computation, we resorted to the double backward trick, which computes the forward-mode gradient with two backward calls (torch.autograd.functional.jvp() in PyTorch).
+This is of course expected to be slower and more memory-intensive, but we can use it as a baseline to verify the validity of the finite-difference approximation employed in the paper.
+For completeness, we also ablated the choice of the step size $\delta$.
+We repeated the box inpainting task of Table 4 and present the results in Table 6.
+In our ablations we find that the finite-difference approximation (i) is robust to the choice of $\delta$, and only fails when using too small (0.0005) and too large (0.5) values, and (ii) performs as well as the exact forward mode auto-differentiation while requiring only a fraction of the time.
+Table 6: Ablation study on the choice of $\delta$ and exact forward-mode auto-differentiation.
+$K$ is the number of steps, $\lambda$ the learning rate and $\delta$ the finite difference step size.
+By * we denote the parameters used for the experiment in the main paper.
+Inpaint (Box)
+Parameters PSNR ↑ LPIPS ↓ FID ↓ Time
+$K=5$, $\lambda=0.5$, $\delta=0.0005$ 17.47 0.37 69.02 15s
+$K=5$, $\lambda=0.5$, $\delta=0.005*$ 18.30 0.30 42.01 15s
+$K=5$, $\lambda=0.5$, $\delta=0.05$ 18.32 0.31 42.44 15s
+$K=5$, $\lambda=0.5$, $\delta=0.5$ 15.80 0.34 61.40 15s
+$K=5$, $\lambda=0.5$, exact 18.27 0.31 44.90 76s
+A.6 DDIM and other sampling methods
+To apply the proposed method, we modified the DDIM sampling algorithm [33].
+We provide a side-by-side comparison to show the difference between the original DDIM and the proposed algorithm.
+Extending our algorithm to other sampling methods should be intuitive by alternating between gradient updates from our algorithm and the diffusion updates computed with the diffusion sampling algorithm used.
+In Algorithms 2, 3 we sketch out a pseudo-algorithm for applying the proposed algorithm to any diffusion solver.
+Beyond DDIM, we also implemented our method with the PNDM scheduler [19] where we get results indistinguishable from DDIM.
+Both DDIM and PNDM implementations are provided in the GitHub repository.
+In Appendix B.4, where we applied our method on rectified flows, we use the Euler ODE solver.
+Again, our method is intuitive to apply by interleaving the diffusion updates with our proposed optimization steps.
+Algorithm 2 Pseudo-algorithm for sampling using a solver and a pre-trained diffusion model.
+1:Input: Pre-trained diffusion model $\hat{{\bm{x}}}_{0}({\bm{x}}_{t})$, diffusion $\text{Solver}()$, diffusion steps $t_{1},t_{2},\dots,t_{N}$, diffusion schedule $\alpha_{i}$
+[…]
+3:for $t=t_{1},t_{2},\dots,t_{N-1}$ do
+4:  ${\bm{z}}_{t}\sim N(\textbf{0},{\bm{I}})$
+5:  ${\bm{x}}_{t+1}=\text{Solver}(\hat{{\bm{x}}}_{0}({\bm{x}}_{t}),{\bm{z}}_{t},\alpha_{t},t+1)$
+6:end for
+7:Return: ${\bm{x}}_{N}$
+Algorithm 3 Pseudo-algorithm for constrained sampling with a solver, a pre-trained diffusion model and using the proposed algorithm.
+1:Input: Pre-trained diffusion model $\hat{{\bm{x}}}_{0}({\bm{x}}_{t})$, diffusion $\text{Solver}()$, constraint $C({\bm{x}}_{0},{\bm{y}})=\lVert f(\hat{{\bm{x}}}_{0}({\bm{x}}_{t})-y\rVert_{2}^{2}$, condition $\bm{y}$}, step size $\delta$, iterations $K$, learning rate $\lambda$, diffusion steps $t_{1},t_{2},\dots,t_{N}$, diffusion schedule $\alpha_{i}$
+[…]
+3:for $t=t_{1},t_{2},\dots,t_{N-1}$ do
+4:  for $i=1,2,\dots,K$ do
+5:   ${\bm{e}}={\bm{J}}_{f}^{T}(f(\hat{{\bm{x}}}_{0}({\bm{x}}_{t}))-{\bm{y}})$
+6:   ${\bm{e}}_{t}=[\hat{{\bm{x}}}_{0}({\bm{x}}_{t}+\delta{\bm{e}})-\hat{{\bm{x}}}_{0}({\bm{x}}_{t})]/\delta$
+7:   ${\bm{x}}_{t}={\bm{x}}_{t}-\lambda{\bm{e}}_{t}$
+8:  end for
+9:  ${\bm{z}}_{t}\sim N(\textbf{0},{\bm{I}})$
+10:  ${\bm{x}}_{t+1}=\text{Solver}(\hat{{\bm{x}}}_{0}({\bm{x}}_{t}),{\bm{z}}_{t},\alpha_{t},t+1)$
+11:end for
+12:Return: ${\bm{x}}_{N}$
+Appendix B Additional results
+B.1 Mask-guided generation
+For the mask-guided generation experiment we utilized a pre-trained face segmentation model from huggingface https://huggingface.co/jonathandinu/face-parsing.
+In Figure 13 we provide qualitative results of our segmentation mask-guided generation experiment described in the main text.
+The quantitative results are provided in Table 5 in the main text.
+Figure 13: Examples of segmentation-guided generation using our method and MPGD.
+B.2 Number of steps and learning rate ablation
 We repeat the ImageNet box inpainting experiments with fewer optimization iterations and a higher learning rate.
@@
 Using a higher learning rate leads to the model sometimes ’overshooting’ by inpainting the missing regions with realistic-looking parts that do not necessarily fit the rest of the image.
-The quantitative results are presented in Table 5.
-Method PSNR ↑ LPIPS ↓ FID ↓
-Ours (n=5, lr=0.5) 18.30 0.30 42.01
-Ours (n=2, lr=0.5) 18.01 0.39 68.75
-Ours (n=5, lr=1.0) 17.48 0.32 47.20
-Table 5: Ablation of the number of steps and learning rate on box inpainting.
-When using fewer steps, the algorithm does not sufficiently converge.
-When using a very high learning rate, the approach seems to overshoot to non-realistic inpainted parts.
+The quantitative results are presented in Table 4 in the main text.
 Figure 14: Examples of box inpainting when using different optimization steps and learning rates.
-C.2 Non-differentiable constraints
+B.3 Non-differentiable constraints
 There is no restriction on defining the constraint $C$ as long as we can get the direction $\bm{e}$} towards which we want to push the image ${\bm{x}}_{0}$.
@@
 As a toy example, we generate images with pixel values quantized to be either ’on’ or ’off’ (-1 or 1).
-The constraint first measures whether a pixel value is positive or negative and then sets the error direction ${\bm{e}}$ to $|1-x|$ or $|-1-x|$ for every pixel accordingly.
+The constraint first measures whether a pixel value is positive or negative and then sets the error direction $\bm{e}$} to |1x||1-x| or |1x||-1-x| for every pixel accordingly.
 This is a non-differentiable constraint for which we can easily compute a local gradient that reduces the cost $C$.
@@
 Figure 15: We apply a non-differentiable constraint to generate quantized images. Using the prompt ’a photo of a cat’ we generated binary images of cats.
-C.3 Inference visualizations
-In Figures 11, 12, and 13, we visualize the intermediate steps of the proposed algorithm for the inpainting, super-resolution and style-guided generation tasks respectively.
+B.4 Time-distilled models
+We test whether our method works on distilled models such as rectified flows [20] and consistency models [35].
+These techniques reduce the number of inference steps by distilling from a base diffusion model.
+Our method does not explicitly depend on the number of inference steps used, the type of model or the noise schedule; the only requirement is having a way to estimate the final clean image from the current step, which both rectified flows and consistency models admit.
+Thus, applying to rectified flows and consistency models is intuitive.
+We employ our method to inpaint images using the 2-rectified flow model distilled from Stable Diffusion, InstaFlow [21], using 5 inference steps.
+We observe that although our algorithm still works with a rectified flow and just 5 inference steps, we do not consistently get high-quality samples as we did with Stable Diffusion when using the same hypeparameters and it requires more optimization steps.
+Ultimately, we utilize line search to find the $\lambda$ for every gradient update we perform.
+This increases inference time further, mitigating the gains from using a time-distilled model.
+We hypothesize that using rectified flows (or any distilled model with fewer steps) may be more challenging since the initial noise dictates most of the content in the final image.
+Therefore, the first optimization steps we perform must get sufficiently close to the correct solution.
+When using a diffusion model, we can get away with imperfect optimization steps as there is more room for ’fixing’ the image in later timesteps.
+We show examples of the rectified flow inpainting in Figure 16.
+We also refer the reviewer to Figures 17,18, where we show the intermediate generation steps for Stable Diffusion.
+Figure 16: Using a rectified flow model [21] that generates images in just 5 steps requires more careful optimization steps, as there is less room for errors during generation.
+B.5 Inference visualizations
+In Figures 17, 18, and 19, we visualize the intermediate steps of the proposed algorithm for the inpainting, super-resolution and style-guided generation tasks respectively.
 Our method quickly converges to a plausible image and then further refines it to better satisfy the constraint over the diffusion timesteps.
@@
 Figure 19: Visualization of the intermediate inference steps for the style-guided generation task.
-C.4 Effect of convergence speed on final images
+B.6 Effect of convergence speed on final images
 We ask the question of how does convergence speed affect the quality of the generated images?
-Previous works found that applying a high weight on the constraint led to unwanted artifacts in the generated images (Chung et al., 2023).
+Previous works found that applying a high weight on the constraint led to unwanted artifacts in the generated images [5].
 We hypothesize that, apart from artifacts in the gradient, ’over-optimizing’ for the condition at a given timestep can affect the generation quality.
@@
 Even when running the optimization for more steps at a single $\mathbf{x}_{t}$, we see that although the sample converges faster to the desired condition, the denoiser is still able to continue the diffusion process of $\mathbf{x}_{t}$.
-Figure 14: Convergence for the inpainting task when using 5 optimization steps $K$.
-Figure 15: Convergence for the inpainting task when using 20 optimization steps $K$.
-C.5 More Qualitative Results
+Figure 20: Convergence for the inpainting task when using $K=5$ optimization steps.
+Figure 21: Convergence for the inpainting task when using $K=20$ optimization steps.
+B.7 More Qualitative Results
 In Figure 3 we provided qualitative results on free-form inpainting and super-resolution.
@@
 In Figure 22 we showcase additional results on the box inpainting task.
-MPGD (He et al., 2024), which does not backpropagate the constraint error through the diffusion model completely fails at inpainting the missing region.
+MPGD [13], which does not backpropagate the constraint error through the diffusion model completely fails at inpainting the missing region.
 We attribute that to the minimal ability to influence pixels that are ’far’ from the constraint at lower noise levels without probing the model weights.
@@
 In Figure 24 we show images generated with different styles and text prompts.
-Here, we show how increasing the classifier-free guidance weight $w$ (Ho & Salimans, 2022) controls the influence of the text prompt on the final generated image.
+Here, we show how increasing the classifier-free guidance weight $w$ [15] controls the influence of the text prompt on the final generated image.
 Figure 23: Examples of style-guided text-to-image generation for a single prompt and multiple styles.
 Figure 24: Examples of style-guided text-to-image generation for multiple prompts and styles.
-In Figure 19 we provide qualitative results of our segmentation mask-guided generation experiment.
-We see that the baseline of MPGD either over-satisfies the constraint by burning in artifacts ($\rho=1$) or fails to generate images that adhere to the constraint ($\rho=0.5$).
-Our approach generates the most realistic images while also getting the mask prediction to match to the reference image.
-Since we are using Stable Diffusion, we used the text prompt ’a headshot photo’ to constrain the set of possible generated images.
-As expected, the results are not on par with an FFHQ-specific model, but we still find our algorithm able to synthesize good images even in this difficult case.
-Figure 19: Examples of segmentation-guided generation using our method and MPGD.
-In Figure 20 we include more samples from the experiment in Section 3.2 in the main paper.
-We observe that the proposed gradient updates regularly give more consistent outputs that repeat textures across the image.
-We attribute this to the difference between ${\bm{J}}$ and ${\bm{J}}^{T}$ we discussed in A.2.
-Figure 20: Comparison between gradient descent and the proposed gradient. The setup is the same one as in Section 3.2.
+Appendix C Societal Impact
+The work presented in this paper aims to advance the field of machine learning, specifically generative modeling.
+Solving constrained sampling tasks with a generative prior, can greatly benefit from the better utilization of the image prior.
+One specific domain is compressed sensing in medical imaging, where generative priors like diffusion models have been used to reconstruct low-dose CT scans and accelerated MRIs.
+We leave to future work the application of the proposed algorithm to these settings.
+However, we acknowledge that there are potential societal consequences of our work that can have a negative impact.
+The one we highlight is the ability to edit images with the intent to deceive and mislead.
+While it is true that existing models can already be used to alter images, we understand that our work could offer more precise control over the generation and lead to more convincingly fabricated content.